### v1

# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control*

Abstract
Practical control systems pose significant challenges in identifying optimal control policies due to uncertainties in the system model and external disturbances. While $H_{\infty}$ control techniques are commonly used to design robust controllers that mitigate the effects of disturbances, these methods often require complex and computationally intensive calculations. To address this issue, this paper proposes a reinforcement learning algorithm called Robust Deterministic Policy Gradient (RDPG), which formulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game. In this formulation, one player (the user) aims to minimize the cost, while the other player (the adversary) seeks to maximize it. We then employ deterministic policy gradient (DPG) and its deep reinforcement learning counterpart to train a robust control policy with effective disturbance attenuation. In particular, for practical implementation, we introduce an algorithm called robust deep deterministic policy gradient (RDDPG), which employs a deep neural network architecture and integrates techniques from the twin-delayed deep deterministic policy gradient (TD3) to enhance stability and learning efficiency.
To evaluate the proposed algorithm, we implement it on an unmanned aerial vehicle (UAV) tasked with following a predefined path in a disturbance-prone environment. The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
I INTRODUCTION
In recent years, significant advancements have been made in deep reinforcement learning (DRL), which integrates deep learning and reinforcement learning (RL). The successes of DRL have been demonstrated across various domains, including games [12, 13] and control [5, 14, 15]. Among the various challenging domains, this paper focuses on dynamic system control problems with continuous state-action spaces.
Practical dynamic control systems pose significant challenges in identifying optimal control policies due to unstructured uncertainties in the system model and external disturbances. Current state-of-the-art DRL methods for continuous control tasks, such as deep deterministic policy gradient (DDPG) [5], twin delayed DDPG (TD3) [4], soft actor-critic (SAC) [2], and proximal policy optimization (PPO) [3], generally fail to address these issues effectively. For instance, a DRL agent trained in a specific real-world or simulated environment may struggle to maintain optimal performance when deployed in different environments where the patterns of external disturbances changed. Therefore, developing a DRL agent policy that is robust to unseen disturbances is crucial for ensuring reliability in real-world applications.
The main goal of this paper is to introduce the robust deterministic policy gradient (RDPG) algorithm, which enhances the robustness of the DPG algorithm by integrating the concept of the $H_{\infty}$ control problem with a two-player zero-sum dynamic game framework.
The $H_{\infty}$ control method is a robust control technique that designs an optimal controller to minimize the impact of disturbances on its performance. It is well-known that, under certain assumptions, the $H_{\infty}$ control problem can be transformed into a two-player zero-sum dynamic game (or a minmax optimization problem), where the first player, or the controller, tries to minimize a specific cost, while the second player, or the adversary, attempts to maximize it [21]. The minmax optimization problem can be solved by the Hamilton–Jacobi–Isaacs (HJI) equation. If the system is linear with quadratic costs, the solution of HJI equation can be derived through a Riccati equation [22, 23]. However, if the system has nonlinearity or the cost is non-quadratic, the HJI equation is extremely difficult to solve directly. Therefore, some methods have tried to obtain the approximate solution of HJI equations including power series expansion-based method [6], locally linearized mode [7], online learning algorithm with neural nets [10], and relaxed policy iteration algorithm with bounded policy evaluation error [11]. However, these methods are still associated with a high computational cost.
In this paper, we present the so-called robust deterministic policy gradient (RDPG) algorithm, which improves the robustness of DPG algorithm from the two-player zero-sum dynamic game perspectives. In particular, we transform the $H_{\infty}$ problem into a two-player zero-sum dynamic game between the user (controller) and the adversary that represents the disturbance. Then, we propose a version of deterministic policy gradient method to train both the user and adversary, where the user and adversary each learn their own policy to minimize and maximize the objective function, respectively. Consequently, the robust controller can determine the optimal control input, which is robust against disturbances. The proposed RDPG is then extended to a DDPG variant, called RDDPG (robust DDPG), which combines the idea of RDPG with TD3.
Additionally, we aim to develop a robust control strategy for quadrotors, a commonly used unmanned aerial vehicle (UAV) across various applications. Ensuring robust quadrotor control is essential, particularly for tasks such as tracking moving targets or following predefined waypoints in dynamic environments. We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances. Furthermore, it outperforms other DRL-based control approaches by achieving lower cost under severe disturbance conditions.
In conclusion, the main contributions can be summarized as follows:
  1. 1.
    We propose the robust deterministic policy gradient (RDPG) algorithm, which effectively improves the robustness against external disturbances by solving the $H_{\infty}$ problem through the dynamic game perspectives in combination with the actor-critic method.
  2. 2.
    Moreover, we develop a new DDPG algorithm, called RDDPG, which combines the concept of RDPG with TD3. Through numerical simulations of the quadrotor control problems, the effectiveness of RDDPG is demonstrated.
  3. 3.
    Through the developed method, we propose a robust tracking control design method for quadrotor applications. Comprehensive experimental results are given to demonstrate that the proposed approach significantly outperforms other state-of-the-art DRL approaches such as DDPG, PPG, SAC, and TD3.
II Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [6, 7, 8, 9]. However, these methods require precise model linearization and incur a high computational costs due to the need to solve HJI equations. To address this issue, alternative approaches leverage deep reinforcement learning (DRL) to determine the optimal control policy, which does not require solving the nonlinear HJI equation.
Until now, the two-player zero-sum game and $H_{\infty}$ control frameworks have been applied to deep reinforcement learning (DRL) to enhance robustness under various scenarios. For instance,[18] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. In [24], the authors proposed a novel approach for robust locomotion control of quadruped robots against external disturbances. A robust adversarial reinforcement learning strategy was proposed in [19] to address modeling errors and discrepancies between training and testing conditions. Although their underlying philosophy is similar to ours, their approaches are significantly different. The method in [24] solves two separate constrained optimization problems based on the ideas in PPO. The DRL in [18] primarily considers model uncertainties rather than external disturbances and applies the so-called action robust RL in [20], which has environmental structures different from our setting. The authors in [19] propose more general alternating optimization procedures and do not consider deterministic policy gradient.
III Background
III-A $H_{\infty}$ Control
Let us consider the discrete time nonlinear system
$x_{k+1}$ $=f(x_{k},u_{k},w_{k},v_{k})$ (1)
$y_{k}$ $=g(x_{k},u_{k},w_{k})$
Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
$x_{k+1}$ $=f(x_{k},\pi(x_{k}),w_{k},v_{k})$
$y_{k}$ $=g(x_{k},\pi(x_{k}),w_{k})$
Assume that the initial state $x_{0}$ is determined by $x_{0}\sim\rho(\cdot)$, where $\rho$ is the initial state distribution. Defining the stochastic processes ${\bf w}_{0:\infty}:=(w_{0},w_{1},\ldots)$ and ${\bf y}_{0:\infty}:=(y_{0},y_{1},\ldots)$, the system can be seen as a stochastic mapping from ${\bf w}_{0:\infty}$ to ${\bf y}_{0:\infty}$ as follows:
${\bf y}_{0:\infty}\sim T_{\pi}(\cdot|{\bf w}_{0:\infty}).$
where $T_{\pi}$ is the conditional probability of ${\bf y}_{0:\infty}$ given ${\bf w}_{0:\infty}$. Moreover, defining the $L^{2}$ norm for the general stochastic process ${\bf z}_{0:\infty}:=(z_{0},z_{1},\ldots)$ by
$||{{\bf{z}}_{0:\infty}}|{|_{{L^{2}}}}:=\sqrt{\sum\limits_{k=0}^{\infty}{\mathbb{E}}[z_{k}^{T}z_{k}]},$
the $H_{\infty}$ norm of the autonomous system is defined as
$||{T_{\pi}}|{|_{\infty}}:=\mathop{\sup}\limits_{{{\bf{w}}_{0:\infty}}\neq 0}\frac{{{{\left\|{{{\bf{y}}_{0:\infty}}}\right\|}_{{L^{2}}}}}}{{{{\left\|{{{\bf{w}}_{0:\infty}}}\right\|}_{{L^{2}}}}}}$ (2)
The goal of $H_{\infty}$ control is to design a control policy $\pi$ that minimizes the $H_{\infty}$ norm of the system $T_{\pi}$. However, minimizing $||T_{\pi}||_{\infty}$ directly is often difficult in practical implementations. Instead, we typically aim to find an acceptable $\eta>0$ and a controller satisfying $||T_{\pi}||_{\infty}\leq\eta$, which is called suboptimal $H_{\infty}$ control problem. Then, the $H_{\infty}$ control problem can be approximated to the problem of finding a controller $\pi$ that satisfies the constraint
$||{T_{\pi}}||_{\infty}^{2}=\mathop{\sup}\limits_{{{\bf{w}}_{0:\infty}}\neq 0}\frac{{||{{\bf{y}}_{0:\infty}}||_{{L^{2}}}^{2}}}{{||{{\bf{w}}_{0:\infty}}||_{{L^{2}}}^{2}}}\leq{\eta^{2}}$ (3)
Defining
${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(y_{k}^{T}{y_{k}}-{\eta^{2}}w_{k}^{T}{w_{k}})}}\right|\pi}\right]$
the problem can be equivalently written by finding a controller $\pi$ satisfying ${J^{\pi}}\leq 0$. The problem can be solved by
${\pi^{*}}:=\arg\mathop{\min}\limits_{\pi}{J^{\pi}}.$
III-B Two-player zero-sum dynamic game perspective
According to [21], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions. In particular, we consider two decision making agents called the user and the adversary, respectively. They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form ${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$, where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the first agent, and $w_{k}$ is the control input of the second agent. The user’s primary goal entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs. There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game. In the alternating two-player dynamic game, two agents engaged in decision making take turns in selecting actions to maximize and minimize the cost, respectively. On the other hand, in the simultaneous two-player dynamic game, the two agents take actions simultaneously to maximize and minimize the cost. In this paper, we mainly focus on the simultaneous two-player dynamic game.
We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$
$u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}}^{n}$
The corresponding cost function is defined as
$J^{\pi,\mu}:={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})\right|\pi,\mu\right]$ (4)
The goal of the dynamic game is to find a saddle-point pair of equilibrium policies $(\pi,\mu)$ (if exists) for which
${J^{{\pi^{*}},\mu}}\leq{J^{\pi*,\mu*}}\leq{J^{\pi,\mu*}},\quad\forall\pi\in\Pi,\quad\forall\mu\in{\rm M}$
where $\Pi$ denotes the set of all admissible state-feedback policies for the user and ${\rm M}$ denotes the set of all admissible state-feedback policies for the adversary. The above relation equivalently means
${\mu^{*}}={\max_{\mu\in{\rm M}}}{J^{{\pi^{*}},\mu}},\quad{\pi^{*}}={\min_{\pi\in\Pi}}{J^{\pi,{\mu^{*}}}}$
By the min-max theorem, this also implies that
${\max_{\mu\in{\rm M}}}{\min_{\pi\in\Pi}}{J^{\pi,\mu}}={\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
The optimal cost is now defined as
${J^{*}}:={J^{{\pi^{*}},{\mu^{*}}}}$
To show a connection between the dynamic game framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
$c(x,u,w)=g{(x,u,w)^{T}}g(x,u,w)-{\eta^{2}}{w^{T}}w$
where $v\sim p(\cdot)$ and $\gamma=1$. The corresponding cost function is then written as
${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(y_{k}^{T}{y_{k}}-{\eta^{2}}w_{k}^{T}{w_{k}})}}\right|\pi,\mu}\right]$
The user’s optimal policy is
${\pi^{*}}:=\arg{\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
which is structurally very similar to the policy with the $H_{\infty}$ performance in the previous subsection. It is known that under some special conditions such as the linearity, the $H_{\infty}$ suboptimal control policy and the saddle-point policy are equivalent when $J^{*}\leq 0$.
III-C Bellman equations
Let us define the value function
${V^{\pi,\mu}}(x):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})}\right|\pi,\mu,x_{t}=x}\right]$
so that the corresponding cost function is $J^{\pi,\mu}={\mathbb{E}}[V^{\pi,\mu}(x)|x\sim\rho(\cdot)]$. We can also prove that the value function satisfies the Bellman equation
$V^{\pi,\mu}(x)=c(x,\pi(x),\mu(x))+\gamma{\mathbb{E}}[{V^{\pi,\mu}}(x^{\prime})|\pi,\mu]$
where $x^{\prime}$ implies the next state given the current state $x$ and the action taken by $(\pi,\mu)$. Let us define the optimal value function as $V^{*}:=V^{\pi^{*},\mu^{*}}$. Then, the corresponding optimal Bellman equation can be obtained by replacing $(\pi,\mu)$ by $(\pi^{*},\mu^{*})$. Similarly, let us define the so-called Q-function by
${Q^{\pi,\mu}}(x,u,w)$
::=:= ${\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=z,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
which satisfies the Q-Bellman equation
${Q^{\pi,\mu}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{\pi,\mu}}(x^{\prime},u^{\prime},w^{\prime})|\pi,\mu]$
where $x^{\prime}$ implies the next state given the current state $x$ and action pair $(u,w)$, $u^{\prime}$ means the next action of the user given $x^{\prime}$ and under $\pi$, and $w^{\prime}$ implies the next action of the adversary given $x^{\prime}$ and under $\mu$. Defining the optimal Q-function $Q^{*}:=Q^{\pi^{*},\mu^{*}}$, one can easily prove the optimal Q-Bellman equation
${Q^{*}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
IV Proposed method
In this section, we will present the proposed method.
IV-A Robust deterministic policy gradient
To apply DPG to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{R^{n}}$
The saddle-point problem is then converted to
${\theta^{*}}:=\arg{\min_{\theta}}{\max_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}$
which can be solved using the primal-dual iteration
${\theta_{k+1}}=$ ${\theta_{k}}-{\alpha_{\theta}}{\left.{{\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{{\phi_{k}}}}}}}\right|_{\theta={\theta_{k}}}},$
${\phi_{k+1}}=$ ${\phi_{k}}+\alpha_{\phi}{\left.{{\nabla_{\phi}}{J^{{\pi_{{\theta_{k+1}}}},{\mu_{\phi}}}}}\right|_{\phi={\phi_{k}}}}$
where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to[1], the deterministic gradients can be obtained using the following theorem.
Theorem 1
The deterministic policy gradients for the user and adversary policies are given by
${\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}=E\left[{\left.{{{\left.{{\nabla_{\theta}}{Q^{\pi,{\mu_{\phi}}}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\pi={\pi_{\theta}}}}}\right|s\sim{\rho^{{\pi_{\theta}},{\mu_{\phi}}}}}\right]$
and
${\nabla_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}=E\left[{\left.{{{\left.{{\nabla_{\phi}}{Q^{{\pi_{\theta}},\mu}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\mu={\mu_{\phi}}}}}\right|s\sim{\rho^{{\pi_{\theta}},{\mu_{\phi}}}}}\right]$
respectively.
The proof is a simple extension of the DPG theorem in [1], so it is omitted here. A reinforcement learning counterpart to DPG can be obtained by using the samples of the gradients. Next, we will consider the following per-step cost function:
$c(x,u,w)=\tilde{c}(x,u,w)-{\eta^{2}}{w^{T}}w$
where $\tilde{c}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{p}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the cost function for the user. Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem. This is because the $H_{\infty}$ control problem can be recovered by setting
$\tilde{c}(x,u,w)=g{(x,u,w,v)^{T}}g(x,u,w,v)$
In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance.
Let us assume that we found some approximate solutions $\theta^{*}_{\epsilon},\phi^{*}_{\epsilon}$ such that $J^{\pi_{\theta_{\varepsilon}^{*}},\mu_{\phi_{\varepsilon}^{*}}}\leq 0$. Then, this implies that
$\frac{{\mathbb{E}}\left[{\left.\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}(x_{k},u_{k},w_{k})}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{{\gamma^{k}}}w_{k}^{T}{w_{k}}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
IV-B Implementation based on TD3 (RDDPG)
To implement RDPG for practical high-dimensional tasks, we will apply the techniques from twin-delayed deep deterministic policy gradient (TD3) in [4], which is called robust deep deterministic policy gradient (RDDPG). In particular, we introduce the following networks:
  1. 1.
    The online actor network $\pi_{\theta}$ for the user and the online actor network $\mu_{\phi}$ for the adversary
  2. 2.
    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two actors
  3. 3.
    Two online critic networks $Q_{\psi_{1}}(x,u,w),Q_{\psi_{2}}(x,u,w)$
  4. 4.
    The corresponding target networks $Q_{\psi_{1}^{\prime}}(x,u,w)$, $Q_{\psi_{2}^{\prime}}(x,u,w)$
Now, following [4], the actor and critic networks are trained through the following procedures.
IV-B1 Critic update
The critic is trained by the gradient descent step to the loss
$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},$
$i\in\{1,2\}$
where $B$ is the mini-batch, $|B|$ is the size of the mini-batch, $c$ is the cost incurred at the same time as the state $x$, and $x^{\prime}$ means the next state. Moreover, the target $y$ is defined as
$y=c+\gamma\max_{i\in\{1,2\}}Q_{\psi_{i}^{\prime}}(x^{\prime},\tilde{u},\tilde{w})$
when $x^{\prime}$ is not the terminal state, and
$y=c$
when $x^{\prime}$ is the terminal state (in the episodic environments), where
$\tilde{u}$ $=\pi_{\theta^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{1}},-c,c)\ \ \epsilon_{1}\sim\mathcal{N}(0,\sigma)$
$\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-c,c)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs. The critic’s online parameters $\phi_{1},\phi_{2}$ are updated by the gradient descent step to minimize the loss
${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
where $\alpha_{\rm critic}$ is the step-size.
IV-B2 Actor update
The actor networks for the user and adversary are updated using the sampled deterministic policy gradient
$\theta\leftarrow\theta-{\alpha_{\rm actor}}{\nabla_{\theta}}{L_{\rm actor}}(\theta,\phi;B)$
$\phi\leftarrow\phi+{\beta_{\rm actor}}{\nabla_{\phi}}{L_{\rm actor}}(\theta,\phi;B)$
where $\alpha_{\rm actor}$ and $\beta_{\rm actor}$ are the step-sizes, and
$L_{\rm actor}(\theta,\phi;B)=\frac{1}{{|B|}}\sum\limits_{(x,u,w,r,x^{\prime})\in B}{{Q_{{\psi_{1}}}}(x,{\pi_{\theta}}(x),{\mu_{\phi}}(x))}$
Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles.
IV-B3 Target parameters update
The target parameters are updated through the averaging
$\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime},\quad\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
for the actor target parameters and
${\psi_{i}}^{\prime}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi_{i}}^{\prime},\quad i\in\{1,2\}$
for the critic target parameters.
IV-B4 Exploration
For exploration, we can use
$u_{k}=$ $\pi_{\theta}(x_{k})+e_{1}$
$w_{k}=$ $\mu_{\phi}(x_{k})+e_{2}$
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [4].
IV-B5 Optimizing $\eta$
As an objective function estimate, we use the target $y$ used above. In particular, we simply assume
${J^{{\pi_{\theta^{\prime}}},{\mu_{\phi^{\prime}}}}}\cong\hat{J}:=\frac{1}{{|B|}}\sum\limits_{(x,u,w,r,x^{\prime})\in B}y$
where the target $y$ is defined as before. Then, we can apply
$J\leftarrow$ $(1-{\alpha_{J}})J+{\alpha_{J}}\hat{J}$
$\eta\leftarrow$ $\eta+\alpha_{\eta}J$
where $J$ is an estimate of the cost function $J^{{\pi_{\theta}},{\mu_{\phi}}}$. The basic idea of the update rule for $\eta$ is as follows: if $J>0$, then increase $\eta>0$ so that $J\leq 0$; if $J<0$, then decrease $\eta>0$ so that $J\leq 0$. In this way, we can find an optimal $\eta$ such that $J\leq 0$. The overall algorithm is summarized in Algorithm 1.
Algorithm 1 RDDPG
1:Initialize the online critic networks $Q_{\psi_{1}},Q_{\psi_{2}}$
2:Initialize the actor networks $\pi_{\theta},\mu_{\phi}$ for the user and adversary, respectively.
3:Initialize the target parameters $\psi_{1}^{\prime}\leftarrow\psi_{1},\psi_{2}^{\prime}\leftarrow\psi_{2}$, $\theta^{\prime}\leftarrow\theta,\phi^{\prime}\leftarrow\phi$
4:Initialize the replay buffer $\mathcal{D}$
5:for  Episode $i$=1,2,…$N_{iter}$ do
6:     Observe $s_{0}$
7:     for  Time step $k$=0,1,2,…$\tau$-1 do
8:         Select actions $u_{k}=\pi_{\theta}(x_{k})+e_{1}$ and $w_{k}=\mu_{\phi}(x_{k})+e_{2}$, where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
9:         Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
10:         Observe the next state $x_{k+1}$
11:         Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
12:         Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
13:         Update critic network:
${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
14:         Update actor networks by the deterministic policy gradient:
$\theta\leftarrow\theta-{\alpha_{\rm actor}}{\nabla_{\theta}}{L_{\rm actor}}(\theta,\phi;B)$
$\phi\leftarrow\phi+{\beta_{\rm actor}}{\nabla_{\phi}}{L_{\rm actor}}(\theta,\phi;B)$
15:         Update $J$ and $\eta$:
$J\leftarrow(1-{\alpha_{J}})J+{\alpha_{J}}\hat{J}$
$\eta\leftarrow\eta+\alpha_{\eta}J$
16:         Soft update target networks:
17:         $\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}^{\prime}$ , $\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}^{\prime}$
18:     end for
19:end for
V Quadrotor application
V-A Quadrotor dynamics model
The quadrotor dynamics can be described as follows:
$m\mathbf{a}_{k}$ $=m\begin{bmatrix}0\\ 0\\ -g\end{bmatrix}+\mathbf{R}\mathbf{T}_{k}+\mathbf{F}_{D}+w_{k}$ (5)
$\Delta\omega_{k}$ $=I^{-1}(-\omega_{k-1}\times I\omega_{k-1}+\tau_{B})$ (6)
$\mathbf{v}_{k}$ $=\mathbf{v}_{k-1}+T_{s}\mathbf{a}_{k}$ (7)
$\mathbf{p}_{k}$ $=\mathbf{p}_{k-1}+T_{s}\mathbf{v}_{k}$ (8)
$\omega_{k}$ $=\omega_{k-1}+T_{s}\Delta\omega_{k}$ (9)
$\rho_{k}$ $=\rho_{k-1}+T_{s}\omega_{k}$ (10)
where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame. Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame. Moreover, $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:
$\mathbf{T}_{k}=\sum_{i=1}^{4}T_{k,i}=k_{F}\begin{bmatrix}0\\ 0\\ \sum\text{RPM}_{k,i}^{2}\\ \end{bmatrix}$
where $T_{k,i}$ is the thrust generated by $i$th motor, $k_{F}$ is the thrust coefficient, and $\text{RPM}_{k,i}$ is the speed of $i$th motor at time step $k$. $\mathbf{F}_{D}$ is a drag force, which is generated by motor spinning, defined as
$\mathbf{F}_{D}=-\mathbf{K}_{D}\left(\sum_{i=1}^{4}\text{RPM}_{k,i}\right)\mathbf{v}_{k}$
where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients. Among the four motors, two opposing motors ($1$th and $3$rd) rotate clockwise, while the others ($2$nd and $4$th) rotate counterclockwise. This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}$
$=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 00, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient. The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
V-B State space
The state $x_{k}$ of the quadrotor is defined as
$x_{k}=[\mathbf{e}_{k}^{\text{pos}},\mathbf{e}_{k}^{\text{vel}},\rho_{k},\omega_{k}]^{T}\in\mathbb{R}^{12}$
where
$\mathbf{e}_{k}^{\text{pos}}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
$=[p_{k,x}^{\text{target}}-p_{k,x},\ p_{k,y}^{\text{target}}-p_{k,y},\ p_{k,z}^{\text{target}}-p_{k,z}]^{T}\in\mathbb{R}^{3}$
$\mathbf{e}_{k}^{\text{vel}}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
$=[v_{k,x}^{\text{target}}-v_{k,x},\ v_{k,y}^{\text{target}}-v_{k,y},\ v_{k,z}^{\text{target}}-v_{k,z}]^{T}\in\mathbb{R}^{3}$
$\rho_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
$\omega_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
where $\mathbf{e}_{k}^{\text{pos}}$ is the position error between the target $\mathbf{p}_{k}^{\text{target}}$ and the quadrotor position $\mathbf{p}_{k}$, $\mathbf{e}_{k}^{\text{vel}}$ is the velocity error between the target $\mathbf{v}_{k}^{\text{target}}$ and the quadrotor $\mathbf{v}_{k}$, $\rho_{k}$ is the Euler angle vector, and $\omega$ represents the angular velocity vector of the quadrotor in its body frame. As in the method used in DQN[13], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
V-C Action space
The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},i\in\{1,2,3,4\}$, which is continuous value between 00 and $21713.714$, as follows:
$u_{k}:=[\text{RPM}_{k,1},\text{RPM}_{k,2},\text{RPM}_{k,3},\text{RPM}_{k,4}]^{T}\in\mathbb{R}^{4}.$
The action of the disturbance adversary, $w_{k}$, represents the external forces to the quadrotor along the $x$, $y$, and $z$ axes of the body frame as follows:
$w_{k}:=[w_{k,x},w_{k,y},w_{k,z}]^{T}\in\mathbb{R}^{3}$
V-D Cost function
The cost function $c(x_{k},u_{k},w_{k})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$. The cost function for the user $\tilde{c}(x_{k},u_{k})$ is the sum of the following terms: $\tilde{c}(x_{k},u_{k})=c_{k,p}+c_{k,v}+c_{k,\rho}+c_{l}+c_{tr}$
  1. 1.
    Position error cost $c_{k,p}$ : Position error between the target and quadrotor
    $c_{k,p}=\alpha\times||\mathbf{e}_{k}^{\text{pos}}||_{2}^{2}$
  2. 2.
    Velocity error cost $c_{k,v}$ : Velocity error between the target and quadrotor
    $c_{k,v}=\beta\times||\mathbf{e}_{k}^{\text{vel}}||_{2}^{2}$
  3. 3.
    Angle cost $c_{k,\rho}$ : Magnitude of roll, pitch, and yaw of the quadrotor
    $c_{k,\rho}=\epsilon\times||\rho_{k}||_{2}^{2}$
  4. 4.
    Living cost $c_{l}$ : Constant cost at every step
    $c_{l}=\lambda$
  5. 5.
    Penalty cost $c_{tr}$ : Penalty when the quadrotor fails to track the target
    $c_{tr}=\begin{cases}\zeta&\text{if }||\mathbf{e}_{k}^{\text{pos}}||_{2}\geq 5\text{ or }\phi_{k},\theta_{k}\geq\frac{\pi}{2}\\ 0&\text{else}\end{cases}$
where $\alpha$, $\beta$, $\epsilon$, $\zeta>0$, and $\lambda<0$ are scaling factors. Since $\alpha$ and $\beta$ are positive, the agent learns a policy that minimizes position and velocity errors. If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily. Therefore, minimizing the angles cost $c_{\rho}$ is crucial for stability. Additionally, the living cost $c_{l}$ is a negative reward that encourages the quadrotor to maintain stability. The penalty cost $c_{tr}$ is a large positive value applied when the quadrotor fails to track the target—specifically, when the distance between the target and the quadrotor exceeds $5$ meters, or when the quadrotor flips with roll or pitch angles exceeding $\frac{\pi}{2}$. The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
VI Experiments and results
VI-A Experiments setups
Figure 1: Trajectories for evaluation.
For experiments, we used gym-pybullet-drones environment [17], which is an open-source quadrotor simulator, built using Python and the Bullet Physics engine. This environment provides a modular and precise physics implementation, supporting both low-level and high-level controls. Crazyflie $2.0$ quadrotor is modeled in the simulator, which is used for our experiments. In experiment, we compare RDDPG with DDPG[5], TD3[4], SAC[2], and PPO[3]. These methods utilize a single agent aiming to minimize the cost function $\tilde{c}$. To enhance their robustness against disturbances, random external disturbances were applied during the learning process. From our own experiences, learning under severe disturbances with the four existing methods failed to yield policies that demonstrated reasonable performance. Therefore, we applied only minor disturbances, limited to ±$\pm 5$ [m/s] when applying the four methods.
For evaluation, we generate three trajectories as shown in Fig. 1. During evaluation, we introduce random wind disturbances in the $x$ and $y$ directions of the inertia frame. We consider four scenarios where the maximum wind velocity is limited to ±$\pm 5$, ±$\pm 10$, ±$\pm 15$, and ±$\pm 20$ [m/s], which are referred to as the ±$\pm 5$, ±$\pm 10$, ±$\pm 15$, and ±$\pm 20$ wind speed scenarios in the tables. In each scenario, the initial wind speed is randomly set within a given range and updated every $10$ time steps with normally distributed noise. Each algorithm was tested over 500 episodes and evaluated using three metrics: the mean and standard deviation of the total cost per episode, the average $H_{\infty}$ norm, empirically evaluated as $\frac{\tilde{c}(x_{k},u_{k},w_{k})}{w_{k}^{T}w_{k}}$, and the mean and standard deviation of the position error between the quadrotor and the target. All metrics were averaged over 500 episodes.
TABLE I: The mean and standard deviation of the cost
Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $\mathbf{-9.20\pm 0.03}$ $\mathbf{-9.11\pm 0.02}$ $\mathbf{-8.94\pm 0.03}$ DDPG $-7.63\pm 0.09$ $-7.67\pm 0.10$ $-7.49\pm 0.12$ TD3 $-4.25\pm 2.47$ $-7.50\pm 1.65$ $-6.00\pm 2.53$ SAC $-7.42\pm 0.98$ $-8.68\pm 0.26$ $-8.34\pm 0.30$ PPO $19.26\pm 12.01$ $7.72\pm 8.37$ $11.03\pm 9.54$$\pm 10$ wind speed scenario RDDPG $\mathbf{-9.00\pm 0.12}$ $\mathbf{-8.96\pm 0.16}$ $\mathbf{-8.73\pm 0.21}$ DDPG $-7.41\pm 0.24$ $-7.41\pm 0.28$ $-7.20\pm 0.43$ TD3 $0.45\pm 5.95$ $-2.20\pm 5.00$ $-0.94\pm 5.00$ SAC $-7.16\pm 1.05$ $-7.94\pm 0.71$ $-7.54\pm 0.87$ PPO $22.89\pm 15.31$ $22.22\pm 15.42$ $22.71\pm 15.64$$\pm 15$ wind speed scenario RDDPG $\mathbf{-8.49\pm 0.52}$ $\mathbf{-8.36\pm 0.71}$ $\mathbf{-8.17\pm 0.62}$ DDPG $-5.95\pm 2.03$ $-5.98\pm 1.90$ $-5.76\pm 1.78$ TD3 $3.02\pm 6.49$ $1.90\pm 6.28$ $2.64\pm 5.45$ SAC $-2.86\pm 4.55$ $-3.70\pm 4.03$ $-3.32\pm 4.13$ PPO $23.63\pm 17.10$ $23.45\pm 16.95$ $23.27\pm 15.77$$\pm 20$ wind speed scenario RDDPG $\mathbf{-5.04\pm 3.90}$ $\mathbf{-4.77\pm 4.07}$ $\mathbf{-4.46\pm 4.03}$ DDPG $0.18\pm 5.56$ $-0.04\pm 5.76$ $-0.51\pm 5.28$ TD3 $3.76\pm 5.32$ $3.27\pm 4.88$ $3.66\pm 4.77$ SAC $0.18\pm 4.64$ $0.49\pm 4.96$ $0.61\pm 4.98$ PPO $20.81\pm 15.58$ $21.15\pm 15.12$ $21.32\pm 16.77$
Figure 2: The cost curve of RDDPG during training. This shows that RDDPG success to minimize the cost as training progresses.
VI-B Results
Fig. 2 shows the sum of costs during the training steps, demonstrating that RDDPG effectively minimizes the cost as training progresses
Table I presents the mean and standard deviation of the total cost sum per episode. Across all scenarios, RDDPG achieved the lowest total cost sum, which demonstrates superior performance compared to other algorithms. Notably, TD3, the baseline algorithm for RDDPG, exhibited significant performance degradation as wind conditions intensified. In contrast, RDDPG maintained a lower cost sum and greater stability, indicating that RDPG enhances the robustness of the DPG algorithm.
Table II describes the mean of $H_{\infty}$ norm at time step $k$ in all wind speed scenarios. As the wind speed decreases, the magnitude of disturbances also diminishes. Since the $H_{\infty}$ norm is inversely proportional to the disturbance magnitude, scenarios with weaker wind conditions tend to exhibit higher $H_{\infty}$ norm values. RDDPG exhibits a smaller $H_{\infty}$ norm than the other algorithms; this implies that RDDPG effectively minimizes the cost in disturbance scenarios
Table III presents the mean and standard deviation of the positional errors between the target and the quadrotor at time step $k$. DDPG performs best in low wind speed scenarios, RDDPG achieves lower errors as wind speed increases. Although DDPG exhibits smaller distance errors, its higher total cost in Table  I suggests that the quadrotor experiences large angular deviations and significant oscillations. These oscillations make the DDPG controller more vulnerable to strong winds, ultimately causing instability in high-wind scenarios.
TABLE II: Empirically evaluated $H_{\infty}$ norm.
Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $\mathbf{247.14}$ $\mathbf{284.04}$ $\mathbf{299.59}$ DDPG $7874.49$ $570.36$ $999.50$ TD3 $1421.04$ $573.83$ $820.78$ SAC $708.79$ $749.96$ $593.65$ PPO $8360.81$ $4086.99$ $7548.96$$\pm 10$ wind speed scenario RDDPG $\mathbf{74.17}$ $\mathbf{112.76}$ $\mathbf{84.36}$ DDPG $185.82$ $282.43$ $202.59$ TD3 $422.65$ $331.02$ $351.93$ SAC $209.82$ $144.25$ $123.55$ PPO $2314.39$ $4058.31$ $2308.93$$\pm 15$ wind speed scenario RDDPG $\mathbf{59.29}$ $\mathbf{44.00}$ $\mathbf{41.09}$ DDPG $75.45$ $87.31$ $82.17$ TD3 $216.96$ $147.81$ $216.73$ SAC $103.45$ $75.54$ $89.90$ PPO $6286.04$ $1281.48$ $2862.71$$\pm 20$ wind speed scenario RDDPG $\mathbf{140.33}$ $\mathbf{30.06}$ $\mathbf{53.50}$ DDPG $76.90$ $83.34$ $79.91$ TD3 $284.40$ $205.07$ $176.43$ SAC $88.42$ $73.83$ $78.32$ PPO $1278.23$ $885.98$ $1576.79$
TABLE III: The mean and standard deviation of the position error $||\mathbf{e}_{k}^{pos}||_{2}$ between the target and the quadrotor
Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $0.17\pm 0.02$ $0.16\pm 0.02$ $0.17\pm 0.03$ DDPG $\mathbf{0.13\pm 0.05}$ $\mathbf{0.12\pm 0.05}$ $\mathbf{0.12\pm 0.07}$ TD3 $0.35\pm 0.28$ $0.21\pm 0.11$ $0.26\pm 0.17$ SAC $0.27\pm 0.16$ $0.20\pm 0.07$ $0.22\pm 0.09$ PPO $0.91\pm 0.65$ $0.86\pm 0.43$ $0.76\pm 0.51$$\pm 10$ wind speed scenario RDDPG ${0.18\pm 0.04}$ ${0.18\pm 0.03}$ ${0.19\pm 0.05}$ DDPG $\mathbf{0.16\pm 0.06}$ $\mathbf{0.15\pm 0.07}$ $\mathbf{0.15\pm 0.08}$ TD3 $0.48\pm 0.39$ $0.38\pm 0.32$ $0.42\pm 0.33$ SAC $0.29\pm 0.16$ $0.25\pm 0.11$ $0.27\pm 0.13$ PPO $1.16\pm 0.82$ $1.13\pm 0.78$ $1.14\pm 0.77$$\pm 15$ wind speed scenario RDDPG $\mathbf{0.22\pm 0.07}$ $\mathbf{0.21\pm 0.07}$ $\mathbf{0.22\pm 0.07}$ DDPG $0.23\pm 0.15$ $0.23\pm 0.15$ $0.23\pm 0.15$ TD3 $0.56\pm 0.42$ $0.52\pm 0.37$ $0.55\pm 0.40$ SAC $0.46\pm 0.28$ $0.43\pm 0.26$ $0.43\pm 0.27$ PPO $1.50\pm 1.16$ $1.50\pm 1.12$ $1.50\pm 1.14$$\pm 20$ wind speed scenario RDDPG $\mathbf{0.22\pm 0.07}$ $\mathbf{0.21\pm 0.07}$ $\mathbf{0.22\pm 0.07}$ DDPG $0.37\pm 0.36$ $0.36\pm 0.36$ $0.35\pm 0.33$ TD3 $0.60\pm 0.47$ $0.57\pm 0.44$ $0.60\pm 0.46$ SAC $0.51\pm 0.36$ $0.51\pm 0.37$ $0.51\pm 0.38$ PPO $1.60\pm 1.28$ $1.57\pm 1.26$ $1.53\pm 1.20$
VII Conclusion
This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user aims to minimize the cost function while the adversary tries to maximize it. Additionally, we introduce the RDDPG algorithm, which integrates RDPG with TD3 to improve both robustness and learning efficiency. In order to evaluate the robustness of RDDPG, we implement it on the tracking tasks of the quadrotor. Experimental results show that RDDPG can achieve the optimal control policy and maintain stability in various environments with disturbances. This demonstrates that RDDPG outperforms other DRL methods in terms of robustness.

### v3

@@
 $y_{k}$ $=g(x_{k},u_{k},w_{k})$
+where $x_{k}\in\mathbb{R}^{n}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{p}$ is the disturbance, and $v_{k}\in\mathbb{R}^{q}$ is the process noise at time step $k$.
 Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
@@
 To apply DPG to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
-$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{R^{n}}$
+$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}^{n}}$
 The saddle-point problem is then converted to
@@
 The deterministic policy gradients for the user and adversary policies are given by
-${\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}=E\left[{\left.{{{\left.{{\nabla_{\theta}}{Q^{\pi,{\mu_{\phi}}}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\pi={\pi_{\theta}}}}}\right|s\sim{\rho^{{\pi_{\theta}},{\mu_{\phi}}}}}\right]$
+${\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}={\mathbb{E}}\left[{\left.{{{\left.{{\nabla_{\theta}}{Q^{\pi,{\mu_{\phi}}}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\pi={\pi_{\theta}}}}}\right|s\sim\rho}\right]$
 and
-${\nabla_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}=E\left[{\left.{{{\left.{{\nabla_{\phi}}{Q^{{\pi_{\theta}},\mu}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\mu={\mu_{\phi}}}}}\right|s\sim{\rho^{{\pi_{\theta}},{\mu_{\phi}}}}}\right]$
+${\nabla_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}={\mathbb{E}}\left[{\left.{{{\left.{{\nabla_{\phi}}{Q^{{\pi_{\theta}},\mu}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\mu={\mu_{\phi}}}}}\right|s\sim\rho}\right]$
 respectively.
@@
 where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame.
-Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame.
-Moreover, $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:
+Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame, and $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:
 $\mathbf{T}_{k}=\sum_{i=1}^{4}T_{k,i}=k_{F}\begin{bmatrix}0\\ 0\\ \sum\text{RPM}_{k,i}^{2}\\ \end{bmatrix}$

### v4

@@
-Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control*
+Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control
 Abstract
@@
 The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
-I INTRODUCTION
-In recent years, significant advancements have been made in deep reinforcement learning (DRL), which integrates deep learning and reinforcement learning (RL).
-The successes of DRL have been demonstrated across various domains, including games [12, 13] and control [5, 14, 15].
-Among the various challenging domains, this paper focuses on dynamic system control problems with continuous state-action spaces.
-Practical dynamic control systems pose significant challenges in identifying optimal control policies due to unstructured uncertainties in the system model and external disturbances.
-Current state-of-the-art DRL methods for continuous control tasks, such as deep deterministic policy gradient (DDPG) [5], twin delayed DDPG (TD3) [4], soft actor-critic (SAC) [2], and proximal policy optimization (PPO) [3], generally fail to address these issues effectively.
-For instance, a DRL agent trained in a specific real-world or simulated environment may struggle to maintain optimal performance when deployed in different environments where the patterns of external disturbances changed.
-Therefore, developing a DRL agent policy that is robust to unseen disturbances is crucial for ensuring reliability in real-world applications.
-The main goal of this paper is to introduce the robust deterministic policy gradient (RDPG) algorithm, which enhances the robustness of the DPG algorithm by integrating the concept of the $H_{\infty}$ control problem with a two-player zero-sum dynamic game framework.
-The $H_{\infty}$ control method is a robust control technique that designs an optimal controller to minimize the impact of disturbances on its performance.
-It is well-known that, under certain assumptions, the $H_{\infty}$ control problem can be transformed into a two-player zero-sum dynamic game (or a minmax optimization problem), where the first player, or the controller, tries to minimize a specific cost, while the second player, or the adversary, attempts to maximize it [21].
-The minmax optimization problem can be solved by the Hamilton–Jacobi–Isaacs (HJI) equation.
-If the system is linear with quadratic costs, the solution of HJI equation can be derived through a Riccati equation [22, 23].
-However, if the system has nonlinearity or the cost is non-quadratic, the HJI equation is extremely difficult to solve directly.
-Therefore, some methods have tried to obtain the approximate solution of HJI equations including power series expansion-based method [6], locally linearized mode [7], online learning algorithm with neural nets [10], and relaxed policy iteration algorithm with bounded policy evaluation error [11].
-However, these methods are still associated with a high computational cost.
-In this paper, we present the so-called robust deterministic policy gradient (RDPG) algorithm, which improves the robustness of DPG algorithm from the two-player zero-sum dynamic game perspectives.
-In particular, we transform the $H_{\infty}$ problem into a two-player zero-sum dynamic game between the user (controller) and the adversary that represents the disturbance.
-Then, we propose a version of deterministic policy gradient method to train both the user and adversary, where the user and adversary each learn their own policy to minimize and maximize the objective function, respectively.
-Consequently, the robust controller can determine the optimal control input, which is robust against disturbances.
-The proposed RDPG is then extended to a DDPG variant, called RDDPG (robust DDPG), which combines the idea of RDPG with TD3.
-Additionally, we aim to develop a robust control strategy for quadrotors, a commonly used unmanned aerial vehicle (UAV) across various applications.
-Ensuring robust quadrotor control is essential, particularly for tasks such as tracking moving targets or following predefined waypoints in dynamic environments.
+keywords:
+Deep reinforcement learning, Two players zero-sum game, Unmanned aerial vehicles
+1 Introduction
+In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4].
+In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models.
+However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment.
+Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications.
+To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances.
+By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
+In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of DPG algorithm from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives.
+The $H_{\infty}$ control method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance.
+A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem.
+In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [8].
+Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between a user agent (controller) and an adversarial agent (disturbance generator).
+We then propose a reinforcement learning approach based on the deterministic policy gradient (DPG) method to simultaneously train both agents.
+Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it.
+As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances.
+Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework.
+The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3.
+Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications.
+The ability to maintain robust control of quadrotors is particularly important in dynamic and uncertain environments, where tasks such as tracking moving targets or following predefined waypoints are required.
 We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances.
-Furthermore, it outperforms other DRL-based control approaches by achieving lower cost under severe disturbance conditions.
+Furthermore, it outperforms existing DRL-based control approaches by achieving lower cost under severe disturbance conditions.
 In conclusion, the main contributions can be summarized as follows:
   1. 1.
-    We propose the robust deterministic policy gradient (RDPG) algorithm, which effectively improves the robustness against external disturbances by solving the $H_{\infty}$ problem through the dynamic game perspectives in combination with the actor-critic method.
+    We propose the robust deterministic policy gradient (RDPG) algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
   2. 2.
-    Moreover, we develop a new DDPG algorithm, called RDDPG, which combines the concept of RDPG with TD3.
-Through numerical simulations of the quadrotor control problems, the effectiveness of RDDPG is demonstrated.
+    We develop a novel deep reinforcement learning algorithm, named RDDPG, which integrates RDPG into the TD3 framework to enable scalable and robust learning in continuous control tasks.
   3. 3.
-    Through the developed method, we propose a robust tracking control design method for quadrotor applications.
-Comprehensive experimental results are given to demonstrate that the proposed approach significantly outperforms other state-of-the-art DRL approaches such as DDPG, PPG, SAC, and TD3.
-II Related works
-The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [6, 7, 8, 9].
-However, these methods require precise model linearization and incur a high computational costs due to the need to solve HJI equations.
-To address this issue, alternative approaches leverage deep reinforcement learning (DRL) to determine the optimal control policy, which does not require solving the nonlinear HJI equation.
-Until now, the two-player zero-sum game and $H_{\infty}$ control frameworks have been applied to deep reinforcement learning (DRL) to enhance robustness under various scenarios.
+    We apply the proposed RDDPG algorithm to the robust tracking control of quadrotors and conduct comprehensive numerical simulations.
+The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL methods, including DDPG, SAC, and TD3.
+2 Related works
+The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [9, 10, 11].
+However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations.
+To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
+Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios.
 For instance, [12] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks.
-In [24], the authors proposed a novel approach for robust locomotion control of quadruped robots against external disturbances.
-A robust adversarial reinforcement learning strategy was proposed in [19] to address modeling errors and discrepancies between training and testing conditions.
-Although their underlying philosophy is similar to ours, their approaches are significantly different.
-The method in [24] solves two separate constrained optimization problems based on the ideas in PPO.
-The DRL in [18] primarily considers model uncertainties rather than external disturbances and applies the so-called action robust RL in [20], which has environmental structures different from our setting.
-The authors in [19] propose more general alternating optimization procedures and do not consider deterministic policy gradient.
-III Background
-III-A $H_{\infty}$ Control
-Let us consider the discrete time nonlinear system
+Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [13], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward.
+In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary.
+By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise.
+In [14], the authors proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator.
+The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning.
+To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [15] was proposed.
+RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system.
+The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it.
+Although the underlying concept is similar, their approach differs significantly from ours.
+Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training.
+In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
+3 Preliminaries
+3.1 $H_{\infty}$ Control
+Let us consider the discrete time nonlinear discrete-time system
 $x_{k+1}$ $=f(x_{k},u_{k},w_{k},v_{k})$ (1)
@@
 Assume that the initial state $x_{0}$ is determined by $x_{0}\sim\rho(\cdot)$, where $\rho$ is the initial state distribution.
-Defining the stochastic processes ${\bf w}_{0:\infty}:=(w_{0},w_{1},\ldots)$ and ${\bf y}_{0:\infty}:=(y_{0},y_{1},\ldots)$, the system can be seen as a stochastic mapping from ${\bf w}_{0:\infty}$ to ${\bf y}_{0:\infty}$ as follows:
-${\bf y}_{0:\infty}\sim T_{\pi}(\cdot|{\bf w}_{0:\infty}).$
-where $T_{\pi}$ is the conditional probability of ${\bf y}_{0:\infty}$ given ${\bf w}_{0:\infty}$.
+Defining the stochastic processes ${\bf w}_{0:\infty}:=(w_{0},w_{1},\ldots)$ and ${\bf y}_{0:\infty}:=(y_{0},y_{1},\ldots)$, the system can be seen as a stochastic mapping from ${\bf w}_{0:\infty}$ to ${\bf y}_{0:\infty}$ as follows: ${\bf y}_{0:\infty}\sim T_{\pi}(\cdot|{\bf w}_{0:\infty})$, where $T_{\pi}$ is the conditional probability of ${\bf y}_{0:\infty}$ given ${\bf w}_{0:\infty}$.
 Moreover, defining the $L^{2}$ norm for the general stochastic process ${\bf z}_{0:\infty}:=(z_{0},z_{1},\ldots)$ by
-$||{{\bf{z}}_{0:\infty}}|{|_{{L^{2}}}}:=\sqrt{\sum\limits_{k=0}^{\infty}{\mathbb{E}}[z_{k}^{T}z_{k}]},$
+[…]
 The $H_{\infty}$ norm of the autonomous system is defined as
-$||{T_{\pi}}|{|_{\infty}}:=\mathop{\sup}\limits_{{{\bf{w}}_{0:\infty}}\neq 0}\frac{{{{\left\|{{{\bf{y}}_{0:\infty}}}\right\|}_{{L^{2}}}}}}{{{{\left\|{{{\bf{w}}_{0:\infty}}}\right\|}_{{L^{2}}}}}}$ (2)
+[…]
 The goal of $H_{\infty}$ control is to design a control policy $\pi$ that minimizes the $H_{\infty}$ norm of the system $T_{\pi}$.
-However, minimizing $||T_{\pi}||_{\infty}$ directly is often difficult in practical implementations.
-Instead, we typically aim to find an acceptable $\eta>0$ and a controller satisfying $||T_{\pi}||_{\infty}\leq\eta$, which is called suboptimal $H_{\infty}$ control problem.
+However, minimizing $T_{\pi}||_{\infty}$ directly is often difficult in practical implementations.
+Instead, we typically aim to find an acceptable $\eta>0$ and a controller satisfying $T_{\pi}||_{\infty}\leq\eta$, which is called suboptimal $H_{\infty}$ control problem.
 Then, the $H_{\infty}$ control problem can be approximated to the problem of finding a controller $\pi$ that satisfies the constraint
-$||{T_{\pi}}||_{\infty}^{2}=\mathop{\sup}\limits_{{{\bf{w}}_{0:\infty}}\neq 0}\frac{{||{{\bf{y}}_{0:\infty}}||_{{L^{2}}}^{2}}}{{||{{\bf{w}}_{0:\infty}}||_{{L^{2}}}^{2}}}\leq{\eta^{2}}$ (3)
+[…]
 Defining
-${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(y_{k}^{T}{y_{k}}-{\eta^{2}}w_{k}^{T}{w_{k}})}}\right|\pi}\right]$
+${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$
 the problem can be equivalently written by finding a controller $\pi$ satisfying ${J^{\pi}}\leq 0$. The problem can be solved by
 ${\pi^{*}}:=\arg\mathop{\min}\limits_{\pi}{J^{\pi}}.$
-III-B Two-player zero-sum dynamic game perspective
+3.2 Two-player zero-sum dynamic game perspective
 According to [8], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions.
 In particular, we consider two decision making agents called the user and the adversary, respectively.
-They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form ${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$, where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the first agent, and $w_{k}$ is the control input of the second agent.
-The user’s primary goal entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs.
+They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form
+${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$
+where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the user, and $w_{k}$ is the control input of the adversary.
+The two agents have opposing objectives.
+The user’s primary objective entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs.
 There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game.
@@
 In this paper, we mainly focus on the simultaneous two-player dynamic game.
-We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$
+We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$ for the adversary
 $u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}}^{n}$
@@
 To show a connection between the dynamic game framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
-$c(x,u,w)=g{(x,u,w)^{T}}g(x,u,w)-{\eta^{2}}{w^{T}}w$
-where $v\sim p(\cdot)$ and $\gamma=1$. The corresponding cost function is then written as
-${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(y_{k}^{T}{y_{k}}-{\eta^{2}}w_{k}^{T}{w_{k}})}}\right|\pi,\mu}\right]$
+$c(x,u,w)=\left\|{g(x,u,w)}\right\|_{2}^{2}-{\eta^{2}}\left\|w\right\|_{2}^{2}.$
+Then the corresponding cost function is then written when $\gamma=1$ as
+${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right]$
 The user’s optimal policy is
@@
 It is known that under some special conditions such as the linearity, the $H_{\infty}$ suboptimal control policy and the saddle-point policy are equivalent when $J^{*}\leq 0$.
-III-C Bellman equations
+3.3 Bellman equations
 Let us define the value function
@@
 Similarly, let us define the so-called Q-function by
-${Q^{\pi,\mu}}(x,u,w)$
-::=:= ${\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=z,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
+${Q^{\pi,\mu}}(x,u,w):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=x,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
 which satisfies the Q-Bellman equation
@@
 ${Q^{*}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
-IV Proposed method
-In this section, we will present the proposed method.
-IV-A Robust deterministic policy gradient
+4 Proposed method
+In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG).
+Figure  1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game.
+In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment.
+The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$.
+These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment.
+The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary.
+The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$.
+The detailed procedure is described in the following sections.
+Figure 1: Overview of RDDPG.
+Based on the two-player zero-sum dynamic game, the user policy $\pi$ seeks to minimize the objective function $J^{\pi,\mu}$, while the adversary policy $\mu$ attempts to maximize it by injecting disturbances into the environment.
+4.1 Robust deterministic policy gradient
 To apply DPG to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
 $u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}^{n}}$
+where $\pi_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ denotes the adversary’s control policy parameterized by $\phi$..
 The saddle-point problem is then converted to
@@
 where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to[16], the deterministic gradients can be obtained using the following theorem.
-Theorem 1
+Theorem 1.
 The deterministic policy gradients for the user and adversary policies are given by
@@
 Next, we will consider the following per-step cost function:
-$c(x,u,w)=\tilde{c}(x,u,w)-{\eta^{2}}{w^{T}}w$
+$c(x,u,w)=\tilde{c}(x,u,w)-\eta^{2}\left\|w\right\|_{2}^{2}$
 where $\tilde{c}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{p}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the cost function for the user.
@@
 This is because the $H_{\infty}$ control problem can be recovered by setting
-$\tilde{c}(x,u,w)=g{(x,u,w,v)^{T}}g(x,u,w,v)$
+$\tilde{c}(x,u,w)=\left\|{g(x,u,w,v)}\right\|_{2}^{2}$
 In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance.
+The additional disturbance quadratic cost ${\eta^{2}}w^{T}w$ in the cost function discourages the adversary from generating excessively large disturbances, effectively regularizing its behavior and ensuring that the learned policies are robust against realistic worst-case scenarios rather than arbitrary or unbounded inputs.
 Let us assume that we found some approximate solutions $\theta^{*}_{\epsilon},\phi^{*}_{\epsilon}$ such that $J^{\pi_{\theta_{\varepsilon}^{*}},\mu_{\phi_{\varepsilon}^{*}}}\leq 0$.
 Then, this implies that
-$\frac{{\mathbb{E}}\left[{\left.\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}(x_{k},u_{k},w_{k})}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{{\gamma^{k}}}w_{k}^{T}{w_{k}}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
-IV-B Implementation based on TD3 (RDDPG)
-To implement RDPG for practical high-dimensional tasks, we will apply the techniques from twin-delayed deep deterministic policy gradient (TD3) in [4], which is called robust deep deterministic policy gradient (RDDPG).
+$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
+4.2 Robust deep deterministic policy gradient
+To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the robust deep deterministic policy gradient (RDDPG).
 In particular, we introduce the following networks:
@@
   2. 2.
-    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two actors
+    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two online actor network
   3. 3.
@@
 Now, following [5], the actor and critic networks are trained through the following procedures.
-IV-B1 Critic update
+4.2.1 Critic update
 The critic is trained by the gradient descent step to the loss
-$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},$
-$i\in\{1,2\}$
+$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},\quad i\in\{1,2\}$
 where $B$ is the mini-batch, $|B|$ is the size of the mini-batch, $c$ is the cost incurred at the same time as the state $x$, and $x^{\prime}$ means the next state.
 Moreover, the target $y$ is defined as
-$y=c+\gamma\max_{i\in\{1,2\}}Q_{\psi_{i}^{\prime}}(x^{\prime},\tilde{u},\tilde{w})$
+[…]
 when $x^{\prime}$ is not the terminal state, and
@@
 when $x^{\prime}$ is the terminal state (in the episodic environments), where
-$\tilde{u}$ $=\pi_{\theta^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{1}},-c,c)\ \ \epsilon_{1}\sim\mathcal{N}(0,\sigma)$
-$\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-c,c)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
+$\tilde{u}$ $=\pi_{\theta^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{1}},-\delta,\delta)\ \ \epsilon_{1}\sim\mathcal{N}(0,\sigma)$
+$\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-\delta,\delta)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
 where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs.
-The critic’s online parameters $\phi_{1},\phi_{2}$ are updated by the gradient descent step to minimize the loss
-${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
+The critic’s online parameters $\psi_{1},\psi_{2}$ are updated by the gradient descent step to minimize the loss
+${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{\nabla_{\psi_{i}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
 where $\alpha_{\rm critic}$ is the step-size.
-IV-B2 Actor update
+4.2.2 Actor update
 The actor networks for the user and adversary are updated using the sampled deterministic policy gradient
@@
 Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles.
-IV-B3 Target parameters update
+4.2.3 Target parameters update
 The target parameters are updated through the averaging
@@
 for the critic target parameters.
-IV-B4 Exploration
+4.2.4 Exploration
 For exploration, we can use
@@
 $w_{k}=$ $\mu_{\phi}(x_{k})+e_{2}$
-where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [4].
-IV-B5 Optimizing $\eta$
-As an objective function estimate, we use the target $y$ used above. In particular, we simply assume
-${J^{{\pi_{\theta^{\prime}}},{\mu_{\phi^{\prime}}}}}\cong\hat{J}:=\frac{1}{{|B|}}\sum\limits_{(x,u,w,r,x^{\prime})\in B}y$
-where the target $y$ is defined as before. Then, we can apply
-$J\leftarrow$ $(1-{\alpha_{J}})J+{\alpha_{J}}\hat{J}$
-$\eta\leftarrow$ $\eta+\alpha_{\eta}J$
-where $J$ is an estimate of the cost function $J^{{\pi_{\theta}},{\mu_{\phi}}}$.
-The basic idea of the update rule for $\eta$ is as follows: if $J>0$, then increase $\eta>0$ so that $J\leq 0$; if $J<0$, then decrease $\eta>0$ so that $J\leq 0$.
-In this way, we can find an optimal $\eta$ such that $J\leq 0$.
-The overall algorithm is summarized in Algorithm 1.
+where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [5]. The overall algorithm is summarized in Algorithm 1.
 Algorithm 1 RDDPG
@@
 4:Initialize the replay buffer $\mathcal{D}$
-5:for  Episode $i$=1,2,…$N_{iter}$ do
+[…]
 6:  Observe $s_{0}$
-7:     for  Time step $k$=0,1,2,…$\tau$-1 do
-8:         Select actions $u_{k}=\pi_{\theta}(x_{k})+e_{1}$ and $w_{k}=\mu_{\phi}(x_{k})+e_{2}$, where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
-9:         Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
-10:         Observe the next state $x_{k+1}$
-11:         Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
-12:         Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
-13:         Update critic network:
+[…]
+8:   Select actions $u_{k}=\pi_{\theta}(x_{k})+e_{1}$ and $w_{k}=\mu_{\phi}(x_{k})+e_{2}$,
+9:   where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
+10:   Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
+11:   Observe the next state $x_{k+1}$
+12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
+13:   Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
+14:   Update critic network:
 ${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
-14:         Update actor networks by the deterministic policy gradient:
+15:   Update actor networks by the deterministic policy gradient:
 $\theta\leftarrow\theta-{\alpha_{\rm actor}}{\nabla_{\theta}}{L_{\rm actor}}(\theta,\phi;B)$
 $\phi\leftarrow\phi+{\beta_{\rm actor}}{\nabla_{\phi}}{L_{\rm actor}}(\theta,\phi;B)$
-15:         Update $J$ and $\eta$:
-$J\leftarrow(1-{\alpha_{J}})J+{\alpha_{J}}\hat{J}$
-$\eta\leftarrow\eta+\alpha_{\eta}J$
 16:   Soft update target networks:
-17:         $\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}^{\prime}$ , $\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}^{\prime}$
+$\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
+$\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}$
+17:  end for
 18:end for
-19:end for
-V Quadrotor application
-V-A Quadrotor dynamics model
+5 Quadrotor application
+5.1 Quadrotor dynamics model
 The quadrotor dynamics can be described as follows:
@@
 where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients.
-Among the four motors, two opposing motors ($1$th and $3$rd) rotate clockwise, while the others ($2$nd and $4$th) rotate counterclockwise.
+Among the four motors, two opposing motors ($1$th and 33rd) rotate clockwise, while the others (22nd and $4$th) rotate counterclockwise.
 This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
-$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}$
-$=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
-where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 00, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient.
+$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
+where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 0, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient.
 The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
-V-B State space
+5.2 State space
 The state $x_{k}$ of the quadrotor is defined as
-$x_{k}=[\mathbf{e}_{k}^{\text{pos}},\mathbf{e}_{k}^{\text{vel}},\rho_{k},\omega_{k}]^{T}\in\mathbb{R}^{12}$
-where
-$\mathbf{e}_{k}^{\text{pos}}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
+$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{\rho}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{12}$
+where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{\rho}_{k}$ is the Euler angle vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector.
+These quantities are defined as follows:
+$\mathbf{e}_{k}^{p}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
 $=[p_{k,x}^{\text{target}}-p_{k,x},\ p_{k,y}^{\text{target}}-p_{k,y},\ p_{k,z}^{\text{target}}-p_{k,z}]^{T}\in\mathbb{R}^{3}$
-$\mathbf{e}_{k}^{\text{vel}}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
+$\mathbf{e}_{k}^{v}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
 $=[v_{k,x}^{\text{target}}-v_{k,x},\ v_{k,y}^{\text{target}}-v_{k,y},\ v_{k,z}^{\text{target}}-v_{k,z}]^{T}\in\mathbb{R}^{3}$
-$\rho_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
-$\omega_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
-where $\mathbf{e}_{k}^{\text{pos}}$ is the position error between the target $\mathbf{p}_{k}^{\text{target}}$ and the quadrotor position $\mathbf{p}_{k}$, $\mathbf{e}_{k}^{\text{vel}}$ is the velocity error between the target $\mathbf{v}_{k}^{\text{target}}$ and the quadrotor $\mathbf{v}_{k}$, $\rho_{k}$ is the Euler angle vector, and $\omega$ represents the angular velocity vector of the quadrotor in its body frame.
+$\boldsymbol{\rho}_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
+$\boldsymbol{\omega}_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
+where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors.
+$\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor.
 As in the method used in DQN[2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
-V-C Action space
-The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},i\in\{1,2,3,4\}$, which is continuous value between 00 and $21713.714$, as follows:
+5.3 Action space
+The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
 $u_{k}:=[\text{RPM}_{k,1},\text{RPM}_{k,2},\text{RPM}_{k,3},\text{RPM}_{k,4}]^{T}\in\mathbb{R}^{4}.$
@@
 $w_{k}:=[w_{k,x},w_{k,y},w_{k,z}]^{T}\in\mathbb{R}^{3}$
-V-D Cost function
+5.4 Cost function
 The cost function $c(x_{k},u_{k},w_{k})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$.
@@
     Position error cost $c_{k,p}$ : Position error between the target and quadrotor
-    $c_{k,p}=\alpha\times||\mathbf{e}_{k}^{\text{pos}}||_{2}^{2}$
+[…]
   2. 2.
     Velocity error cost $c_{k,v}$ : Velocity error between the target and quadrotor
-    $c_{k,v}=\beta\times||\mathbf{e}_{k}^{\text{vel}}||_{2}^{2}$
+[…]
   3. 3.
     Angle cost $c_{k,\rho}$ : Magnitude of roll, pitch, and yaw of the quadrotor
-    $c_{k,\rho}=\epsilon\times||\rho_{k}||_{2}^{2}$
+[…]
   4. 4.
@@
     Penalty cost $c_{tr}$ : Penalty when the quadrotor fails to track the target
-    $c_{tr}=\begin{cases}\zeta&\text{if }||\mathbf{e}_{k}^{\text{pos}}||_{2}\geq 5\text{ or }\phi_{k},\theta_{k}\geq\frac{\pi}{2}\\ 0&\text{else}\end{cases}$
-where $\alpha$, $\beta$, $\epsilon$, $\zeta>0$, and $\lambda<0$ are scaling factors.
-Since $\alpha$ and $\beta$ are positive, the agent learns a policy that minimizes position and velocity errors.
+[…]
+where the scaling factors $\alpha$, $\beta$, $\epsilon$, and $\zeta$ are positive (i.e., $\alpha,\beta,\epsilon,\zeta>0$), and $\lambda$ is negative (i.e., $\lambda<0$).
+The specific values used for these parameters are summarized in the Appendix  2.
+Since $\alpha$ and $\beta$ are positive, the quadrotor agent learns a policy that minimizes position and velocity errors.
 If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily.
@@
 The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
-VI Experiments and results
-VI-A Experiments setups
-Figure 1: Trajectories for evaluation.
-For experiments, we used gym-pybullet-drones environment [17], which is an open-source quadrotor simulator, built using Python and the Bullet Physics engine.
-This environment provides a modular and precise physics implementation, supporting both low-level and high-level controls.
-Crazyflie $2.0$ quadrotor is modeled in the simulator, which is used for our experiments.
-In experiment, we compare RDDPG with DDPG[5], TD3[4], SAC[2], and PPO[3].
-These methods utilize a single agent aiming to minimize the cost function $\tilde{c}$.
-To enhance their robustness against disturbances, random external disturbances were applied during the learning process.
-From our own experiences, learning under severe disturbances with the four existing methods failed to yield policies that demonstrated reasonable performance.
-Therefore, we applied only minor disturbances, limited to ±$\pm 5$ [m/s] when applying the four methods.
-For evaluation, we generate three trajectories as shown in Fig.
-1.
-During evaluation, we introduce random wind disturbances in the $x$ and $y$ directions of the inertia frame.
-We consider four scenarios where the maximum wind velocity is limited to ±$\pm 5$, ±$\pm 10$, ±$\pm 15$, and ±$\pm 20$ [m/s], which are referred to as the ±$\pm 5$, ±$\pm 10$, ±$\pm 15$, and ±$\pm 20$ wind speed scenarios in the tables.
-In each scenario, the initial wind speed is randomly set within a given range and updated every $10$ time steps with normally distributed noise.
-Each algorithm was tested over 500 episodes and evaluated using three metrics: the mean and standard deviation of the total cost per episode, the average $H_{\infty}$ norm, empirically evaluated as $\frac{\tilde{c}(x_{k},u_{k},w_{k})}{w_{k}^{T}w_{k}}$, and the mean and standard deviation of the position error between the quadrotor and the target.
-All metrics were averaged over 500 episodes.
-TABLE I: The mean and standard deviation of the cost
-Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $\mathbf{-9.20\pm 0.03}$ $\mathbf{-9.11\pm 0.02}$ $\mathbf{-8.94\pm 0.03}$ DDPG $-7.63\pm 0.09$ $-7.67\pm 0.10$ $-7.49\pm 0.12$ TD3 $-4.25\pm 2.47$ $-7.50\pm 1.65$ $-6.00\pm 2.53$ SAC $-7.42\pm 0.98$ $-8.68\pm 0.26$ $-8.34\pm 0.30$ PPO $19.26\pm 12.01$ $7.72\pm 8.37$ $11.03\pm 9.54$$\pm 10$ wind speed scenario RDDPG $\mathbf{-9.00\pm 0.12}$ $\mathbf{-8.96\pm 0.16}$ $\mathbf{-8.73\pm 0.21}$ DDPG $-7.41\pm 0.24$ $-7.41\pm 0.28$ $-7.20\pm 0.43$ TD3 $0.45\pm 5.95$ $-2.20\pm 5.00$ $-0.94\pm 5.00$ SAC $-7.16\pm 1.05$ $-7.94\pm 0.71$ $-7.54\pm 0.87$ PPO $22.89\pm 15.31$ $22.22\pm 15.42$ $22.71\pm 15.64$$\pm 15$ wind speed scenario RDDPG $\mathbf{-8.49\pm 0.52}$ $\mathbf{-8.36\pm 0.71}$ $\mathbf{-8.17\pm 0.62}$ DDPG $-5.95\pm 2.03$ $-5.98\pm 1.90$ $-5.76\pm 1.78$ TD3 $3.02\pm 6.49$ $1.90\pm 6.28$ $2.64\pm 5.45$ SAC $-2.86\pm 4.55$ $-3.70\pm 4.03$ $-3.32\pm 4.13$ PPO $23.63\pm 17.10$ $23.45\pm 16.95$ $23.27\pm 15.77$$\pm 20$ wind speed scenario RDDPG $\mathbf{-5.04\pm 3.90}$ $\mathbf{-4.77\pm 4.07}$ $\mathbf{-4.46\pm 4.03}$ DDPG $0.18\pm 5.56$ $-0.04\pm 5.76$ $-0.51\pm 5.28$ TD3 $3.76\pm 5.32$ $3.27\pm 4.88$ $3.66\pm 4.77$ SAC $0.18\pm 4.64$ $0.49\pm 4.96$ $0.61\pm 4.98$ PPO $20.81\pm 15.58$ $21.15\pm 15.12$ $21.32\pm 16.77$
-Figure 2: The cost curve of RDDPG during training. This shows that RDDPG success to minimize the cost as training progresses.
-VI-B Results
-Fig. 2 shows the sum of costs during the training steps, demonstrating that RDDPG effectively minimizes the cost as training progresses
-Table I presents the mean and standard deviation of the total cost sum per episode.
-Across all scenarios, RDDPG achieved the lowest total cost sum, which demonstrates superior performance compared to other algorithms.
-Notably, TD3, the baseline algorithm for RDDPG, exhibited significant performance degradation as wind conditions intensified.
-In contrast, RDDPG maintained a lower cost sum and greater stability, indicating that RDPG enhances the robustness of the DPG algorithm.
-Table II describes the mean of $H_{\infty}$ norm at time step $k$ in all wind speed scenarios.
-As the wind speed decreases, the magnitude of disturbances also diminishes.
-Since the $H_{\infty}$ norm is inversely proportional to the disturbance magnitude, scenarios with weaker wind conditions tend to exhibit higher $H_{\infty}$ norm values.
-RDDPG exhibits a smaller $H_{\infty}$ norm than the other algorithms; this implies that RDDPG effectively minimizes the cost in disturbance scenarios
-Table III presents the mean and standard deviation of the positional errors between the target and the quadrotor at time step $k$.
-DDPG performs best in low wind speed scenarios, RDDPG achieves lower errors as wind speed increases.
-Although DDPG exhibits smaller distance errors, its higher total cost in Table  I suggests that the quadrotor experiences large angular deviations and significant oscillations.
-These oscillations make the DDPG controller more vulnerable to strong winds, ultimately causing instability in high-wind scenarios.
-TABLE II: Empirically evaluated $H_{\infty}$ norm.
-Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $\mathbf{247.14}$ $\mathbf{284.04}$ $\mathbf{299.59}$ DDPG $7874.49$ $570.36$ $999.50$ TD3 $1421.04$ $573.83$ $820.78$ SAC $708.79$ $749.96$ $593.65$ PPO $8360.81$ $4086.99$ $7548.96$$\pm 10$ wind speed scenario RDDPG $\mathbf{74.17}$ $\mathbf{112.76}$ $\mathbf{84.36}$ DDPG $185.82$ $282.43$ $202.59$ TD3 $422.65$ $331.02$ $351.93$ SAC $209.82$ $144.25$ $123.55$ PPO $2314.39$ $4058.31$ $2308.93$$\pm 15$ wind speed scenario RDDPG $\mathbf{59.29}$ $\mathbf{44.00}$ $\mathbf{41.09}$ DDPG $75.45$ $87.31$ $82.17$ TD3 $216.96$ $147.81$ $216.73$ SAC $103.45$ $75.54$ $89.90$ PPO $6286.04$ $1281.48$ $2862.71$$\pm 20$ wind speed scenario RDDPG $\mathbf{140.33}$ $\mathbf{30.06}$ $\mathbf{53.50}$ DDPG $76.90$ $83.34$ $79.91$ TD3 $284.40$ $205.07$ $176.43$ SAC $88.42$ $73.83$ $78.32$ PPO $1278.23$ $885.98$ $1576.79$
-TABLE III: The mean and standard deviation of the position error $||\mathbf{e}_{k}^{pos}||_{2}$ between the target and the quadrotor
-Bold indicates the best performance algorithm Algorithm Trajectory (1) (2) (3) ±$\pm 5$ wind speed scenario RDDPG $0.17\pm 0.02$ $0.16\pm 0.02$ $0.17\pm 0.03$ DDPG $\mathbf{0.13\pm 0.05}$ $\mathbf{0.12\pm 0.05}$ $\mathbf{0.12\pm 0.07}$ TD3 $0.35\pm 0.28$ $0.21\pm 0.11$ $0.26\pm 0.17$ SAC $0.27\pm 0.16$ $0.20\pm 0.07$ $0.22\pm 0.09$ PPO $0.91\pm 0.65$ $0.86\pm 0.43$ $0.76\pm 0.51$$\pm 10$ wind speed scenario RDDPG ${0.18\pm 0.04}$ ${0.18\pm 0.03}$ ${0.19\pm 0.05}$ DDPG $\mathbf{0.16\pm 0.06}$ $\mathbf{0.15\pm 0.07}$ $\mathbf{0.15\pm 0.08}$ TD3 $0.48\pm 0.39$ $0.38\pm 0.32$ $0.42\pm 0.33$ SAC $0.29\pm 0.16$ $0.25\pm 0.11$ $0.27\pm 0.13$ PPO $1.16\pm 0.82$ $1.13\pm 0.78$ $1.14\pm 0.77$$\pm 15$ wind speed scenario RDDPG $\mathbf{0.22\pm 0.07}$ $\mathbf{0.21\pm 0.07}$ $\mathbf{0.22\pm 0.07}$ DDPG $0.23\pm 0.15$ $0.23\pm 0.15$ $0.23\pm 0.15$ TD3 $0.56\pm 0.42$ $0.52\pm 0.37$ $0.55\pm 0.40$ SAC $0.46\pm 0.28$ $0.43\pm 0.26$ $0.43\pm 0.27$ PPO $1.50\pm 1.16$ $1.50\pm 1.12$ $1.50\pm 1.14$$\pm 20$ wind speed scenario RDDPG $\mathbf{0.22\pm 0.07}$ $\mathbf{0.21\pm 0.07}$ $\mathbf{0.22\pm 0.07}$ DDPG $0.37\pm 0.36$ $0.36\pm 0.36$ $0.35\pm 0.33$ TD3 $0.60\pm 0.47$ $0.57\pm 0.44$ $0.60\pm 0.46$ SAC $0.51\pm 0.36$ $0.51\pm 0.37$ $0.51\pm 0.38$ PPO $1.60\pm 1.28$ $1.57\pm 1.26$ $1.53\pm 1.20$
-VII Conclusion
-This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms.
-In RDPG, the user aims to minimize the cost function while the adversary tries to maximize it.
-Additionally, we introduce the RDDPG algorithm, which integrates RDPG with TD3 to improve both robustness and learning efficiency.
-In order to evaluate the robustness of RDDPG, we implement it on the tracking tasks of the quadrotor.
-Experimental results show that RDDPG can achieve the optimal control policy and maintain stability in various environments with disturbances.
-This demonstrates that RDDPG outperforms other DRL methods in terms of robustness.
+6 Experiments and results
+6.1 Experiments setups
+For experiments, we used the gym-pybullet-drones environment [17], which is an open-source quadrotor simulator built with Python and the Bullet Physics engine.
+This environment provided a modular and precise physics implementation, supporting both low-level and high-level control.
+In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in Appendix Table 1.
+To demonstrate the robustness of the proposed RDDPG algorithm, we compared it with other model-free reinforcement learning algorithms, including DDPG [3], TD3 [5], and SAC [6].
+These methods utilized a single agent that aimed to minimize the cost function $\tilde{c}$.
+To enhance their robustness against disturbances, random external disturbances were applied during the training process.
+A detailed summary of the hyperparameters used in training was provided in Appendix 3 and  4.
+For evaluation, we generated two trajectories as shown in Fig.
+2.
+During evaluation, we introduced random wind disturbances in the $x$ and $y$ directions.
+We considered three scenarios where the maximum wind velocity was limited to $\pm 5$, $\pm 10$, and $\pm 15$ [m/s], which were referred to as the $\pm 5$, $\pm 10$, and $\pm 15$ wind speed scenarios in the figures and tables.
+In each scenario, the initial wind speed was randomly set within a given range and updated every $10$ time steps with normally distributed noise.
+Each algorithm was tested over 500 episodes and evaluated based on the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k})$ over a single episode.
+In addition, the mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each timestep were also considered as evaluation metrics.
+Traj (1)
+Traj (2)
+Figure 2: Two evaluation trajectories, denoted as Traj (1) and Traj (2), shown from left to right.
+6.2 Results
+Figure 3: Mean and standard deviation of the total episode cost.
+RDDPG achieves the lowest average cost and the smallest variance across all scenarios.
+This indicates that the RDDPG yields a robust agent that performs consistently even under severe external disturbances.
+[…]
+By maintaining a consistently low and stable cost across a wide range of disturbance intensities, RDDPG exhibits superior robustness and reliable control performance, even under challenging environmental conditions.
+This enhanced robustness stems from its adversarial training framework, inspired by $H_{\infty}$ control theory and formulated as a two-player zero-sum game.
+By optimizing the policy against worst-case disturbances during training, RDDPG learns disturbance-resilient strategies that improve generalization and ensure stable performance under unseen or highly uncertain conditions.
+Figure 4: Trajectory results of all algorithms under the $\pm 15$ wind scenario and Traj (1).
+Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds.
+RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms.
+The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure  5.
+Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC.
+In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios.
+However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG.
+Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes.
+Detailed numerical results are provided in at Appendix Table  5.
+Figure 5: Mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each time step.
+RDDPG consistently achieves low error and standard deviation across all scenarios, whereas other methods exhibit unstable performance.
+7 Conclusion
+This paper proposes robust deterministic policy gradient (RDPG), which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms.
+In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances.
+Furthermore, we introduce robust deep deterministic policy gradient (RDDPG), which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks.
+In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances.
+Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios.
+These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.
+Appendix A Hyperparameters Used in Training
+A.1 Physical parameters of Crazyflies 2.0
+Table 1: Physical parameters of the Crazyfile 2.0
+Parameter Description Value
+m Total mass 0.0270.027[kg]
+l Arm length 0.03970.0397[m]
+$K_{F}$ Thrust coefficient $3.16\times 10^{-10}$
+$K_{M}$ Torque coefficient $7.94\times 10^{-12}$
+$I_{xx}$ Principal Moment of Inertia around x axis $1.4\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
+$I_{yy}$ Principal Moment of Inertia around y axis $1.4\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
+$I_{zz}$ Principal Moment of Inertia around z axis $2.17\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
+$K_{D.xy}$ Drag coefficients in $x,y$ axes $9.18\times 10^{-7}$
+$K_{D.z}$ Drag coefficients in $z$ axis $10.31\times 10^{-7}$
+A.2 Cost function
+Table 2: Scaling factors in cost function
+Parameter Description Value
+$\alpha$ Coefficient for position error $10$
+$\beta$ Coefficient for velocity error 11
+$\epsilon$ Coefficient for angle error 11
+$\lambda$ Living reward 10-10
+$\zeta$ Penalty $1000$
+A.3 DDPG and TD3
+Table 3: Caption
+Parameter Description Value
+$\alpha_{actor}$ Learning rate for user 0.0001
+$\beta_{actor}$ Learning rate for adversary 0.0001
+$\alpha_{critic}$ Learning rate for critic 0.001
+$|B|$ Batch size 256
+$\gamma$ Discounted factor 0.99
+$\sigma$ Exploration noise 0.2
+A.4 SAC
+Table 4: SAC hyper-parameters
+Parameter Description Value
+$\alpha_{actor}$ Learning rate for user 0.0003
+$\alpha_{critic}$ Learning rate for critic 0.003
+$|B|$ Batch size 256
+$\gamma$ Discounted factor 0.99
+$\alpha_{entropy}$ Entropy coefficient 0.99
+Appendix B Experiment results
+Table 5: The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}$ between the target and the quadrotor. Bold indicates the best performance algorithm
+$\pm 5$ wind speed $\pm 10$ wind speed $\pm 15$ wind speed
+Traj (1) Traj (2) Traj (1) Traj (2) Traj (1) Traj (2)
+[…]
+[…]
+[…]
+[…]
+Figure 6: In $\pm 15$ m/s wind scenario
+Figure 7: In $\pm 15$ m/s wind scenario
+Figure 8: In $\pm 15$ m/s wind scenario
+Figure 9: In $\pm 15$ m/s wind scenario
+Figure 10: In $\pm 15$ m/s wind scenario

### v5

@@
-Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control
+Robust Deterministic Policy Gradient in Zero-Sum Games for Disturbance Attenuation
 Abstract
@@
 In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models.
-However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment.
+However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment [8, 9, 10].
 Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications.
@@
 By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
-In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of DPG algorithm from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives.
-The $H_{\infty}$ control method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance.
-A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem.
+In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of deterministic policy gradient (DPG) algorithm [11] from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives.
+The $H_{\infty}$ control [12, 13, 14] method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance.
+A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem [15].
 In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [15].
-Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between a user agent (controller) and an adversarial agent (disturbance generator).
-We then propose a reinforcement learning approach based on the deterministic policy gradient (DPG) method to simultaneously train both agents.
+Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between the user (the controller) and the adversarial agent.
+We then propose a reinforcement learning approach based on the DPG method to simultaneously train both agents.
 Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it.
@@
 Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework.
-The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3.
+The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3 [5].
 Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications.
@@
   1. 1.
-    We propose the robust deterministic policy gradient (RDPG) algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
+    We propose the RDPG algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
   2. 2.
@@
 2 Related works
-The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [9, 10, 11].
+The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [14, 16, 17, 18].
 However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations.
@@
 Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios.
-For instance, [12] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks.
+For instance, Deshpande et al.
+[19] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks.
 Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [9], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward.
@@
 By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise.
-In [14], the authors proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator.
+Long et al.
+[10] proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator.
 The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning.
@@
 $y_{k}$ $=g(x_{k},u_{k},w_{k})$
-where $x_{k}\in\mathbb{R}^{n}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{p}$ is the disturbance, and $v_{k}\in\mathbb{R}^{q}$ is the process noise at time step $k$.
+where $x_{k}\in\mathbb{R}^{p}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{n}$ is the disturbance, and $v_{k}\in\mathbb{R}^{l}$ is the process noise at time step $k$.
 Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
@@
 In this paper, we mainly focus on the simultaneous two-player dynamic game.
-We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$ for the adversary
-$u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}}^{n}$
+We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{p}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{p}\to{\mathbb{R}}^{n}$ for the adversary
+$u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{n},\quad x\in{\mathbb{R}}^{p}$
 The corresponding cost function is defined as
@@
 In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG).
-Figure  1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game.
+Figure 1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game.
 In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment.
@@
 4.1 Robust deterministic policy gradient
-To apply DPG to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
-$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}^{n}}$
-where $\pi_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ denotes the adversary’s control policy parameterized by $\phi$..
+To apply deterministic policy gradient (DPG) [11] to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
+$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{n},\quad x\in{\mathbb{R}^{p}}$
+where $\pi_{\theta}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{n}$ denotes the adversary’s control policy parameterized by $\phi$.
 The saddle-point problem is then converted to
@@
 $c(x,u,w)=\tilde{c}(x,u,w)-\eta^{2}\left\|w\right\|_{2}^{2}$
-where $\tilde{c}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{p}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the cost function for the user.
+where $\tilde{c}:{\mathbb{R}}^{p}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\to{\mathbb{R}}$ is the cost function for the user.
 Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem.
@@
 4.2 Robust deep deterministic policy gradient
-To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the robust deep deterministic policy gradient (RDDPG).
+To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the RDDPG.
 In particular, we introduce the following networks:
@@
 ${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{\nabla_{\psi_{i}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
-where $\alpha_{\rm critic}$ is the step-size.
+where $\alpha_{\rm critic}$ is the step-size. After the online critics are updated, the target parameters of two critic $\psi^{\prime}_{1}$ and $\psi^{\prime}_{2}$ are updated as follows:
+${\psi^{\prime}_{i}}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi^{\prime}_{i}},\quad i\in\{1,2\}$
+where $\tau\in(0,1)$ serves as the interpolation coefficient, facilitating a gradual update that enhances training stability.
 4.2.2 Actor update
@@
 Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles.
-4.2.3 Target parameters update
-The target parameters are updated through the averaging
-$\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime},\quad\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
-for the actor target parameters and
-${\psi_{i}}^{\prime}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi_{i}}^{\prime},\quad i\in\{1,2\}$
-for the critic target parameters.
+After the online actors are updated, the target parameters of actor $\theta^{\prime}$ and $\phi^{\prime}$ are updated as follows
+$\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime},\quad\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
+where $\tau\in(0,1)$ is the interpolation coefficient.
 4.2.3 Exploration
-For exploration, we can use
+For exploration during training, Gaussian noise is added to both the user’s control action and the adversary’s disturbance.
+Specifically, the user’s action $u_{k}$ and the adversary’s disturbance $w_{k}$ at time step $k$ are expressed as
 $u_{k}=$ $\pi_{\theta}(x_{k})+e_{1}$
 $w_{k}=$ $\mu_{\phi}(x_{k})+e_{2}$
-where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [5]. The overall algorithm is summarized in Algorithm 1.
+where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ represent zero-mean Gaussian exploration noises with standard deviation $\sigma$ [5]. The overall algorithm is summarized in Algorithm 1.
 Algorithm 1 RDDPG
@@
 9:   where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
-10:   Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
+10:   Compute the cost $c_{k+1}:=c(x_{k},u_{k},w_{k})$
 11:   Observe the next state $x_{k+1}$
-12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
+12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k+1},x_{k+1})$ in the replay buffer $D$
 13:   Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
@@
 16:   Soft update target networks:
-$\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
-$\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}$
+$\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
+$\phi_{i}^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
 17:  end for
@@
 $\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor.
-As in the method used in DQN[2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
+As in the method used in deep Q-network (DQN) [2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
 5.3 Action space
@@
 RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms.
-The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure  5.
+The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure 5.
 Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC.
@@
 Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes.
-Detailed numerical results are provided in at Appendix Table  5.
+Detailed numerical results are provided in at Appendix Table 5.
 Figure 5: Mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each time step.
@@
 7 Conclusion
-This paper proposes robust deterministic policy gradient (RDPG), which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms.
+This paper proposes RDPG, which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms.
 In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances.
-Furthermore, we introduce robust deep deterministic policy gradient (RDDPG), which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks.
+Furthermore, we introduce RDDOG, which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks.
 In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances.

### v6

@@
-Robust Deterministic Policy Gradient in Zero-Sum Games for Disturbance Attenuation
+Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control
 Abstract
-Practical control systems pose significant challenges in identifying optimal control policies due to uncertainties in the system model and external disturbances.
-While $H_{\infty}$ control techniques are commonly used to design robust controllers that mitigate the effects of disturbances, these methods often require complex and computationally intensive calculations.
-To address this issue, this paper proposes a reinforcement learning algorithm called robust deterministic policy gradient (RDPG), which formulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game.
-In this formulation, one player (the user) aims to minimize the cost, while the other player (the adversary) seeks to maximize it.
-We then employ deterministic policy gradient (DPG) and its deep reinforcement learning counterpart to train a robust control policy with effective disturbance attenuation.
-In particular, for practical implementation, we introduce an algorithm called robust deep deterministic policy gradient (RDDPG), which employs a deep neural network architecture and integrates techniques from the twin-delayed deep deterministic policy gradient (TD3) to enhance stability and learning efficiency.
-To evaluate the proposed algorithm, we implement it on an unmanned aerial vehicle (UAV) tasked with following a predefined path in a disturbance-prone environment.
-The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
+This paper presents a robust reinforcement learning algorithm called robust deterministic policy gradient (RDPG), which reformulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game between a user and an adversary.
+The method combines deterministic policy gradients with deep reinforcement learning to train a robust policy that attenuates disturbances efficiently.
+A practical variant, robust deep deterministic policy gradient (RDDPG), integrates twin-delayed updates for stability and sample efficiency.
+Experiments on an unmanned aerial vehicle demonstrate superior robustness and tracking accuracy under severe disturbance conditions.
 keywords:
-Deep reinforcement learning, Two players zero-sum game, Unmanned aerial vehicles
+Reinforcement learning and deep learning in control, Robust control applications, Trajectory tracking and path following for AVs, AI and learning-based control for automotive systems.
+thanks: This work was supported by the Institute of Information Communications Technology Planning Evaluation (IITP) funded by the Korea government under Grant 2022-0-00469.
 1 Introduction
-In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4].
-In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models.
-However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment [8, 9, 10].
-Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications.
-To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances.
-By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
-In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of deterministic policy gradient (DPG) algorithm [11] from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives.
-The $H_{\infty}$ control [12, 13, 14] method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance.
-A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem [15].
-In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [15].
-Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between the user (the controller) and the adversarial agent.
-We then propose a reinforcement learning approach based on the DPG method to simultaneously train both agents.
-Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it.
-As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances.
+In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal control policies across various domains, including games (RL1; DQN) and control tasks (DDPG; RL2), where various model-free algorithms have demonstrated the ability to generate effective control inputs without requiring explicit system models (DDPG; TD3; SAC; PPO).
+However, in real-world applications, physical systems are inevitably subject to external disturbances, unmodeled dynamics, and variations in system parameters.
+These factors often lead to significant discrepancies between the training simulation environments and the conditions encountered in practice (RARL; AR-MDP; HRL).
+Consequently, policies trained in simulation frequently suffer from performance degradation or instability when deployed on real systems.
+This highlights that developing DRL policies that remain robust under unseen disturbances and model uncertainties is essential for ensuring reliability in practical applications.
+In classical control theory, the $H_{\infty}$ control framework (H1; H2; Hinf1) provides a principled approach for designing controllers that minimize the worst-case impact of disturbances on system performance.
+A well-known result shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game (TZDG), or equivalently, a min–max optimization problem (Hbook).
+In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances (Hbook).
+Building upon this theoretical foundation, we introduce the robust deterministic policy gradient (RDPG) algorithm, which enhances the robustness of the deterministic policy gradient (DPG) method (DPG) by transforming the $H_{\infty}$ control formulation into a TZDG between the user (controller) and an adversarial agent.
+We develop a reinforcement learning framework in which both agents are trained simultaneously: the user aims to minimize the objective, while the adversary attempts to maximize it.
+This min–max interplay naturally encourages the user to acquire a policy that is robust to worst-case disturbances.
 Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework.
-The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3 [5].
+The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3 (TD3).
 Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications.
 The ability to maintain robust control of quadrotors is particularly important in dynamic and uncertain environments, where tasks such as tracking moving targets or following predefined waypoints are required.
-We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances.
-Furthermore, it outperforms existing DRL-based control approaches by achieving lower cost under severe disturbance conditions.
+To this end, we demonstrate that the proposed RDDPG method is capable of generating optimal control inputs even in the presence of external disturbances.
+Experimental results show that RDDPG achieves lower costs and exhibits stronger robustness under disturbance conditions compared to existing DRL-based control approaches.
 In conclusion, the main contributions can be summarized as follows:
@@
   3. 3.
-    We apply the proposed RDDPG algorithm to the robust tracking control of quadrotors and conduct comprehensive numerical simulations.
-The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL methods, including DDPG, SAC, and TD3.
+    We apply the proposed RDDPG algorithm to the tracking control of quadrotor and conduct comprehensive numerical simulations.
+The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL and robust RL methods.
 2 Related works
-The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [14, 16, 17, 18].
+The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems (Hinf1; Hinf2; Hinf3; Hinf5).
 However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations.
@@
 Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios.
-For instance, Deshpande et al.
-[19] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks.
-Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [9], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward.
+For instance, RRLforquad developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks.
+Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework (AR-MDP), where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward.
 In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary.
 By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise.
-Long et al.
-[10] proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator.
+HRL proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator.
 The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning.
-To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [8] was proposed.
+To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) (RARL) was proposed.
 RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system.
@@
 Although the underlying concept is similar, their approach differs significantly from ours.
-Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training.
+Both methods adopt stochastic policy frameworks and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training.
 In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
@@
 Let us consider the discrete time nonlinear discrete-time system
-$x_{k+1}$ $=f(x_{k},u_{k},w_{k},v_{k})$ (1)
+$x_{k+1}$ $=f(x_{k},u_{k},w_{k},v_{k})$
 $y_{k}$ $=g(x_{k},u_{k},w_{k})$
@@
 Defining
-${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$
+${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$ (3)
 the problem can be equivalently written by finding a controller $\pi$ satisfying ${J^{\pi}}\leq 0$. The problem can be solved by
 ${\pi^{*}}:=\arg\mathop{\min}\limits_{\pi}{J^{\pi}}.$
-3.2 Two-player zero-sum dynamic game perspective
-According to [15], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions.
-In particular, we consider two decision making agents called the user and the adversary, respectively.
-They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form
-${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$
-where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the user, and $w_{k}$ is the control input of the adversary.
-The two agents have opposing objectives.
-The user’s primary objective entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs.
-There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game.
-In the alternating two-player dynamic game, two agents engaged in decision making take turns in selecting actions to maximize and minimize the cost, respectively.
-On the other hand, in the simultaneous two-player dynamic game, the two agents take actions simultaneously to maximize and minimize the cost.
-In this paper, we mainly focus on the simultaneous two-player dynamic game.
+3.2 Two-player zero-sum dynamic game
+According to (Hbook), the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a two-player zero-sum dynamic game (TZDG) under certain assumptions.
+In particular, we consider two decision making agents called a user and an adversary, respectively.
+TZDG can be expressed as a tuple $(\mathcal{S},\mathcal{A}_{1},\mathcal{A}_{2},P,c,\gamma,x_{0})$.
+Here $\mathcal{S}=\mathbb{R}^{p}$ is the state space, $\mathcal{A}_{1}=\mathbb{R}^{m}$ and $\mathcal{A}_{2}=\mathbb{R}^{n}$ are the continuous action spaces for the user and adversary, respectively.
+The user selects an control input $u\in\mathcal{A}_{1}$ and the second agent selects disturbance $w\in\mathcal{A}_{2}$ at the current state $x\in\mathcal{S}$ simultaneously.
+Then the state transits to the next state $x^{\prime}\in\mathcal{S}$ with the state transition probability $P(x^{\prime}|x,u,w)$ and the cost $c\in\mathbb{R}$ is incurred by the cost function $c(x,u,w,x^{\prime}):\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\times\mathcal{S}\rightarrow\mathbb{R}$.
+The $\gamma\in(0,1]$ is the discounted factor and $x_{0}\in\mathcal{S}$ represents the initial state.
 We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{p}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{p}\to{\mathbb{R}}^{n}$ for the adversary
 $u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{n},\quad x\in{\mathbb{R}}^{p}$
-The corresponding cost function is defined as
-$J^{\pi,\mu}:={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})\right|\pi,\mu\right]$ (4)
+The objective of the user and adversary are to minimize and maximize the cumulative discounted costs over infinite time horizon $J^{\pi,\mu}$ defined as
+$J^{\pi,\mu}:={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})\right|\pi,\mu\right]$
+where $\mathbb{E}[\cdot|\pi,\mu]$ is an expectation conditioned on the two policies $\pi$ and $\mu$.
 The goal of the dynamic game is to find a saddle-point pair of equilibrium policies $(\pi,\mu)$ (if exists) for which
@@
 ${\mu^{*}}={\max_{\mu\in{\rm M}}}{J^{{\pi^{*}},\mu}},\quad{\pi^{*}}={\min_{\pi\in\Pi}}{J^{\pi,{\mu^{*}}}}$
-By the min-max theorem, this also implies that
-${\max_{\mu\in{\rm M}}}{\min_{\pi\in\Pi}}{J^{\pi,\mu}}={\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
-The optimal cost is now defined as
-${J^{*}}:={J^{{\pi^{*}},{\mu^{*}}}}$
-To show a connection between the dynamic game framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
-$c(x,u,w)=\left\|{g(x,u,w)}\right\|_{2}^{2}-{\eta^{2}}\left\|w\right\|_{2}^{2}.$
+perolat2015approximate; AR-MDP demonstrated that for a game with optimal equilibrium return $J^{*}$, there always exists the Nash equilibrium, and it is equivalent to the minimax solution,
+${J^{*}}:={J^{{\pi^{*}},{\mu^{*}}}}={\max_{\mu\in{\rm M}}}{\min_{\pi\in\Pi}}{J^{\pi,\mu}}={\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$ (4)
+To show a connection between this TZDG framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
+$c(x_{k},u_{k},w_{k},x_{k+1})=\left\|{g(x_{k},u_{k},w_{k},x_{k+1})}\right\|_{2}^{2}-{\eta^{2}}\left\|w_{k}\right\|_{2}^{2}.$
 Then the corresponding cost function is then written when $\gamma=1$ as
-${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right]$
+${J^{\pi,\mu}}={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})\right|\pi,\mu\right]$
+$={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|g(x_{k},u_{k},w_{k},x_{k+1})\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right].$ (5)
 The user’s optimal policy is
 ${\pi^{*}}:=\arg{\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
-which is structurally very similar to the policy with the $H_{\infty}$ performance in the previous subsection.
+which is structurally very similar to Equation (3), the policy with the $H_{\infty}$ performance in the previous subsection.
 It is known that under some special conditions such as the linearity, the $H_{\infty}$ suboptimal control policy and the saddle-point policy are equivalent when $J^{*}\leq 0$.
@@
 Let us define the value function
-${V^{\pi,\mu}}(x):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})}\right|\pi,\mu,x_{t}=x}\right]$
+${V^{\pi,\mu}}(x):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})}\right|\pi,\mu,x_{k}=x}\right]$
 so that the corresponding cost function is $J^{\pi,\mu}={\mathbb{E}}[V^{\pi,\mu}(x)|x\sim\rho(\cdot)]$. We can also prove that the value function satisfies the Bellman equation
-$V^{\pi,\mu}(x)=c(x,\pi(x),\mu(x))+\gamma{\mathbb{E}}[{V^{\pi,\mu}}(x^{\prime})|\pi,\mu]$
+$V^{\pi,\mu}(x)=c(x,\pi(x),\mu(x),x^{\prime})+\gamma{\mathbb{E}}[{V^{\pi,\mu}}(x^{\prime})|\pi,\mu]$
 where $x^{\prime}$ implies the next state given the current state $x$ and the action taken by $(\pi,\mu)$.
@@
 Similarly, let us define the so-called Q-function by
-${Q^{\pi,\mu}}(x,u,w):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=x,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
+${Q^{\pi,\mu}}(x,u,w):=$
+${\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})}}\right|{x_{k}}=x,{u_{k}}=u,{w_{k}}=w,\pi,\mu}\right]$
 which satisfies the Q-Bellman equation
-${Q^{\pi,\mu}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{\pi,\mu}}(x^{\prime},u^{\prime},w^{\prime})|\pi,\mu]$
+${Q^{\pi,\mu}}(x,u,w)=c(x,u,w,x^{\prime})+\gamma{\mathbb{E}}[{Q^{\pi,\mu}}(x^{\prime},u^{\prime},w^{\prime})|\pi,\mu]$
 where $x^{\prime}$ implies the next state given the current state $x$ and action pair $(u,w)$, $u^{\prime}$ means the next action of the user given $x^{\prime}$ and under $\pi$, and $w^{\prime}$ implies the next action of the adversary given $x^{\prime}$ and under $\mu$.
 Defining the optimal Q-function $Q^{*}:=Q^{\pi^{*},\mu^{*}}$, one can easily prove the optimal Q-Bellman equation
-${Q^{*}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
+${Q^{*}}(x,u,w)=c(x,u,w,x^{\prime})+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
 4 Proposed method
 In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG).
-Figure 1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game.
-In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment.
-The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$.
-These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment.
-The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary.
-The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$.
+Fig.
+1 illustrates the architecture of the proposed method, RDDPG, based on the TZDG.
+In this method, two policy, user policy $\pi_{\theta}$ and adversary policy $\mu_{\phi}$, interact within the environment.
+The user policy generates the action $u_{k}$, while the adversary policy generates the disturbance $w_{k}$, both based on the current state $x_{k}$.
+These inputs are fed into the environment, which returns the next state $x_{k+1}$ and the cost $c_{k+1}$ based on the dynamics of the environment.
+The critic is updated based on $x_{k},u_{k},w_{k},x_{k+1},c_{k+1}$} and evaluates the joint performance of the user and adversary.
+The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi_{\theta},\mu_{\phi}}$.
 The detailed procedure is described in the following sections.
 Figure 1: Overview of RDDPG.
-Based on the two-player zero-sum dynamic game, the user policy $\pi$ seeks to minimize the objective function $J^{\pi,\mu}$, while the adversary policy $\mu$ attempts to maximize it by injecting disturbances into the environment.
+Based on the two-player zero-sum dynamic game, the user policy $\pi_{\theta}$ seeks to minimize the objective function $J^{\pi_{\theta},\mu_{\phi}}$, while the adversary policy $\mu_{\phi}$ attempts to maximize it by injecting disturbances into the environment.
 4.1 Robust deterministic policy gradient
-To apply deterministic policy gradient (DPG) [11] to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
+To apply deterministic policy gradient (DPG) (DPG) to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
 $u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{n},\quad x\in{\mathbb{R}^{p}}$
@@
 ${\phi_{k+1}}=$ ${\phi_{k}}+\alpha_{\phi}{\left.{{\nabla_{\phi}}{J^{{\pi_{{\theta_{k+1}}}},{\mu_{\phi}}}}}\right|_{\phi={\phi_{k}}}}$
-where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to [11], the deterministic gradients can be obtained using the following theorem.
-Theorem 1.
+where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to DPG, the deterministic gradients can be obtained using the following theorem.
+Theorem 1
 The deterministic policy gradients for the user and adversary policies are given by
@@
 respectively.
-The proof is a simple extension of the DPG theorem in [11], so it is omitted here.
+The proof is a simple extension of the DPG theorem (DPG), so it is omitted here.
 A reinforcement learning counterpart to DPG can be obtained by using the samples of the gradients.
 Next, we will consider the following per-step cost function:
-$c(x,u,w)=\tilde{c}(x,u,w)-\eta^{2}\left\|w\right\|_{2}^{2}$
+$c(x_{k},u_{k},w_{k},x_{k+1})=\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})-\eta^{2}\left\|w_{k}\right\|_{2}^{2}$
 where $\tilde{c}:{\mathbb{R}}^{p}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\to{\mathbb{R}}$ is the cost function for the user.
 Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem.
-This is because the $H_{\infty}$ control problem can be recovered by setting
-$\tilde{c}(x,u,w)=\left\|{g(x,u,w,v)}\right\|_{2}^{2}$
-In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance.
+This is because the $H_{\infty}$ control problem can be recovered as shown Equation (3) and (5) by setting
+$\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})=\left\|{g(x_{k},u_{k},w_{k},x_{k+1})}\right\|_{2}^{2}$
+In this way, we can consider the TZDG as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance.
 The additional disturbance quadratic cost ${\eta^{2}}w^{T}w$ in the cost function discourages the adversary from generating excessively large disturbances, effectively regularizing its behavior and ensuring that the learned policies are robust against realistic worst-case scenarios rather than arbitrary or unbounded inputs.
@@
 Then, this implies that
-$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
+$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}},{x_{k+1}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
 4.2 Robust deep deterministic policy gradient
-To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the RDDPG.
+To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3) (TD3), resulting in the RDDPG.
 In particular, we introduce the following networks:
   1. 1.
-    The online actor network $\pi_{\theta}$ for the user and the online actor network $\mu_{\phi}$ for the adversary
+    The online actor networks $\pi_{\theta}(x)$ and $\mu_{\phi}(x)$ for the user adversary respectively
   2. 2.
-    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two online actor network
+    The corresponding target actor networks $\pi_{\theta^{\prime}}(x)$, $\mu_{\phi^{\prime}}(x)$
   3. 3.
@@
   4. 4.
-    The corresponding target networks $Q_{\psi_{1}^{\prime}}(x,u,w)$, $Q_{\psi_{2}^{\prime}}(x,u,w)$
-Now, following [5], the actor and critic networks are trained through the following procedures.
+    The corresponding target critic networks
+    $Q_{\psi_{1}^{\prime}}(x,u,w)$, $Q_{\psi_{2}^{\prime}}(x,u,w)$
+Now, following TD3, the actor and critic networks are trained through the following procedures.
 4.2.1 Critic update
 The critic is trained by the gradient descent step to the loss
-$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},\quad i\in\{1,2\}$
-where $B$ is the mini-batch, $|B|$ is the size of the mini-batch, $c$ is the cost incurred at the same time as the state $x$, and $x^{\prime}$ means the next state.
-Moreover, the target $y$ is defined as
+$L_{\rm critic}$ $(\psi_{i};B):=$
+$\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},i\in\{1,2\}$
+where $B$ is the mini-batch, $|B|$ is the size of the mini-batch. Moreover, the target $y$ is defined as
 […]
@@
 $\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-\delta,\delta)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
-where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs.
-The critic’s online parameters $\psi_{1},\psi_{2}$ are updated by the gradient descent step to minimize the loss
+where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs with bound $\delta$.
+The critic’s online parameters $\psi_{1}$ and $\psi_{2}$ are updated by the gradient descent step to minimize the loss $L_{\rm critic}(\psi_{i};B)$
 ${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{\nabla_{\psi_{i}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
-where $\alpha_{\rm critic}$ is the step-size. After the online critics are updated, the target parameters of two critic $\psi^{\prime}_{1}$ and $\psi^{\prime}_{2}$ are updated as follows:
-${\psi^{\prime}_{i}}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi^{\prime}_{i}},\quad i\in\{1,2\}$
-where $\tau\in(0,1)$ serves as the interpolation coefficient, facilitating a gradual update that enhances training stability.
 4.2.2 Actor update
@@
 $L_{\rm actor}(\theta,\phi;B)=\frac{1}{{|B|}}\sum\limits_{(x,u,w,r,x^{\prime})\in B}{{Q_{{\psi_{1}}}}(x,{\pi_{\theta}}(x),{\mu_{\phi}}(x))}$
-Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles.
-After the online actors are updated, the target parameters of actor $\theta^{\prime}$ and $\phi^{\prime}$ are updated as follows
-$\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime},\quad\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
-where $\tau\in(0,1)$ is the interpolation coefficient.
-4.2.3 Exploration
-For exploration during training, Gaussian noise is added to both the user’s control action and the adversary’s disturbance.
-Specifically, the user’s action $u_{k}$ and the adversary’s disturbance $w_{k}$ at time step $k$ are expressed as
-$u_{k}=$ $\pi_{\theta}(x_{k})+e_{1}$
-$w_{k}=$ $\mu_{\phi}(x_{k})+e_{2}$
-where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ represent zero-mean Gaussian exploration noises with standard deviation $\sigma$ [5]. The overall algorithm is summarized in Algorithm 1.
+Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite objectives.
 Algorithm 1 RDDPG
@@
 5 Quadrotor application
-5.1 Quadrotor dynamics model
-The quadrotor dynamics can be described as follows:
-$m\mathbf{a}_{k}$ $=m\begin{bmatrix}0\\ 0\\ -g\end{bmatrix}+\mathbf{R}\mathbf{T}_{k}+\mathbf{F}_{D}+w_{k}$ (5)
-$\Delta\omega_{k}$ $=I^{-1}(-\omega_{k-1}\times I\omega_{k-1}+\tau_{B})$ (6)
-$\mathbf{v}_{k}$ $=\mathbf{v}_{k-1}+T_{s}\mathbf{a}_{k}$ (7)
-$\mathbf{p}_{k}$ $=\mathbf{p}_{k-1}+T_{s}\mathbf{v}_{k}$ (8)
-$\omega_{k}$ $=\omega_{k-1}+T_{s}\Delta\omega_{k}$ (9)
-$\rho_{k}$ $=\rho_{k-1}+T_{s}\omega_{k}$ (10)
-where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame.
-Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame, and $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:
-$\mathbf{T}_{k}=\sum_{i=1}^{4}T_{k,i}=k_{F}\begin{bmatrix}0\\ 0\\ \sum\text{RPM}_{k,i}^{2}\\ \end{bmatrix}$
-where $T_{k,i}$ is the thrust generated by $i$th motor, $k_{F}$ is the thrust coefficient, and $\text{RPM}_{k,i}$ is the speed of $i$th motor at time step $k$.
-$\mathbf{F}_{D}$ is a drag force, which is generated by motor spinning, defined as
-$\mathbf{F}_{D}=-\mathbf{K}_{D}\left(\sum_{i=1}^{4}\text{RPM}_{k,i}\right)\mathbf{v}_{k}$
-where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients.
-Among the four motors, two opposing motors ($1$th and 33rd) rotate clockwise, while the others (22nd and $4$th) rotate counterclockwise.
-This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
-$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
-where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 0, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient.
-The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
 5.1 State space
 The state $x_{k}$ of the quadrotor is defined as
-$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{\rho}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{12}$
-where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{\rho}_{k}$ is the Euler angle vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector.
+$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{q}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{13}$
+where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{q}_{k}$ is the quaternion vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector.
 These quantities are defined as follows:
 $\mathbf{e}_{k}^{p}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
-$=[p_{k,x}^{\text{target}}-p_{k,x},\ p_{k,y}^{\text{target}}-p_{k,y},\ p_{k,z}^{\text{target}}-p_{k,z}]^{T}\in\mathbb{R}^{3}$
+$=[p_{x,k}^{\text{target}}-p_{x,k},\ p_{y,k}^{\text{target}}-p_{y,k},\ p_{z,k}^{\text{target}}-p_{z,k}]^{T}\in\mathbb{R}^{3}$
 $\mathbf{e}_{k}^{v}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
-$=[v_{k,x}^{\text{target}}-v_{k,x},\ v_{k,y}^{\text{target}}-v_{k,y},\ v_{k,z}^{\text{target}}-v_{k,z}]^{T}\in\mathbb{R}^{3}$
-$\boldsymbol{\rho}_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
-$\boldsymbol{\omega}_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
+$=[v_{x,k}^{\text{target}}-v_{x,k},\ v_{k,y}^{\text{target}}-v_{y,k},\ v_{z,k}^{\text{target}}-v_{z,k}]^{T}\in\mathbb{R}^{3}$
+$\boldsymbol{q}_{k}$ $=[q_{w,k},q_{x,k},q_{y,k}q_{z,k}]^{T}\in\mathbb{R}^{4}$
+$\boldsymbol{\omega}_{k}$ $=[\omega_{\phi,k},\omega_{\theta,k},\omega_{\psi,k}]^{T}\in\mathbb{R}^{3}$
 where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors.
-$\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor.
-As in the method used in deep Q-network (DQN) [2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
+As in the method used in deep Q-network (DQN) (DQN), we used the last four time steps of the state as input to the neural networks to help the network better capture dynamics in the environment.
 5.2 Action space
-The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
-$u_{k}:=[\text{RPM}_{k,1},\text{RPM}_{k,2},\text{RPM}_{k,3},\text{RPM}_{k,4}]^{T}\in\mathbb{R}^{4}.$
-The action of the disturbance adversary, $w_{k}$, represents the external forces to the quadrotor along the $x$, $y$, and $z$ axes of the body frame as follows:
-$w_{k}:=[w_{k,x},w_{k,y},w_{k,z}]^{T}\in\mathbb{R}^{3}$
+The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{i,k},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
+$u_{k}:=[\text{RPM}_{1,k},\text{RPM}_{2,k},\text{RPM}_{3,k},\text{RPM}_{4,k}]^{T}\in\mathbb{R}^{4}.$
+The disturbance of the adversary, $w_{k}$, represents the external forces to the quadrotor along the $x$, $y$, and $z$ axes of the body frame of the quadrotor as follows:
+$w_{k}:=[w_{x,k},w_{y,k},w_{z,k}]^{T}\in\mathbb{R}^{3}$
+For training stability, each component of the disturbance was constrained to lie within the range [0.1,,0.1][-0.1,,0.1].
 5.3 Cost function
-The cost function $c(x_{k},u_{k},w_{k})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$.
-The cost function for the user $\tilde{c}(x_{k},u_{k})$ is the sum of the following terms: $\tilde{c}(x_{k},u_{k})=c_{k,p}+c_{k,v}+c_{k,\rho}+c_{l}+c_{tr}$
+The cost function $c(x_{k},u_{k},w_{k},x_{k+1})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$.
+The cost function for the user $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})$ is the sum of the following terms: $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})=c_{p,k}+c_{v,k}+c_{\rho,k}+c_{tr}$
   1. 1.
-    Position error cost $c_{k,p}$ : Position error between the target and quadrotor
+    Position error cost $c_{p,k}$ : Position error between the target and quadrotor
 […]
   2. 2.
-    Velocity error cost $c_{k,v}$ : Velocity error between the target and quadrotor
+    Velocity error cost $c_{v,k}$ : Velocity error between the target and quadrotor
 […]
   3. 3.
-    Angle cost $c_{k,\rho}$ : Magnitude of roll, pitch, and yaw of the quadrotor
+    Angle cost $c_{\rho,k}$ : Magnitude of roll $\phi_{k}$, pitch $\theta_{k}$, and yaw $\psi_{k}$ of the quadrotor
 […]
   4. 4.
-    Living cost $c_{l}$ : Constant cost at every step
-    $c_{l}=\lambda$
-  5. 5.
     Penalty cost $c_{tr}$ : Penalty when the quadrotor fails to track the target
@@
 where the scaling factors $\alpha$, $\beta$, $\epsilon$, and $\zeta$ are positive (i.e., $\alpha,\beta,\epsilon,\zeta>0$), and $\lambda$ is negative (i.e., $\lambda<0$).
-The specific values used for these parameters are summarized in the Appendix  2.
-Since $\alpha$ and $\beta$ are positive, the quadrotor agent learns a policy that minimizes position and velocity errors.
-If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily.
-Therefore, minimizing the angles cost $c_{\rho}$ is crucial for stability.
-Additionally, the living cost $c_{l}$ is a negative reward that encourages the quadrotor to maintain stability.
-The penalty cost $c_{tr}$ is a large positive value applied when the quadrotor fails to track the target—specifically, when the distance between the target and the quadrotor exceeds $5$ meters, or when the quadrotor flips with roll or pitch angles exceeding $\frac{\pi}{2}$.
-The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
+The specific values used for these parameters are summarized in the Appendix of arxiv version.
 6 Experiments and results
 6.1 Experiments setups
-For experiments, we used the gym-pybullet-drones environment [20], which is an open-source quadrotor simulator built with Python and the Bullet Physics engine.
-This environment provided a modular and precise physics implementation, supporting both low-level and high-level control.
-In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in Appendix Table 1.
-To demonstrate the robustness of the proposed RDDPG algorithm, we compared it with other model-free reinforcement learning algorithms, including DDPG [3], TD3 [5], and SAC [6].
-These methods utilized a single agent that aimed to minimize the cost function $\tilde{c}$.
-To enhance their robustness against disturbances, random external disturbances were applied during the training process.
-A detailed summary of the hyperparameters used in training was provided in Appendix 3 and  4.
-For evaluation, we generated two trajectories as shown in Fig.
-2.
-During evaluation, we introduced random wind disturbances in the $x$ and $y$ directions.
-We considered three scenarios where the maximum wind velocity was limited to $\pm 5$, $\pm 10$, and $\pm 15$ [m/s], which were referred to as the $\pm 5$, $\pm 10$, and $\pm 15$ wind speed scenarios in the figures and tables.
-In each scenario, the initial wind speed was randomly set within a given range and updated every $10$ time steps with normally distributed noise.
-Each algorithm was tested over 500 episodes and evaluated based on the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k})$ over a single episode.
-In addition, the mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each timestep were also considered as evaluation metrics.
+Figure 2: Mean and standard deviation of the total episode cost.
+RDDPG achieves the lowest average cost and the smallest variance.
+This indicates that the RDDPG yields a robust agent that performs consistently even under external disturbances.
 Traj (1)
@@
 Figure 3: Two evaluation trajectories, denoted as Traj (1) and Traj (2), shown from left to right.
+For experiments, we used the gym-pybullet-drones environment (gym), which is an open-source quadrotor simulator built with Python and the Bullet Physics engine.
+This environment provided a modular and precise physics implementation, supporting both low-level and high-level control.
+In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in the accompanying arXiv version of the paper.
+To evaluate the robustness of the proposed RDDPG algorithm, we compare it against deterministic policy gradient–based reinforcement learning methods, including TD3 (TD3), which employs a single agent that seeks to minimize the cost function $\tilde{c}$.
+Beyond these standard single-agent baseline, we further benchmark RDDPG against robust reinforcement learning algorithms such as RARL (RARL), PR-MDP, and NR-MDP (AR-MDP).
+These robust methods are formulated under the TZDG framework, where the user and adversary aim to minimize and maximize the cost function, respectively.
+The full set of hyperparameters used in our experiments is provided in the accompanying arXiv version of the paper.
+For evaluation, we generated two trajectories as shown in Fig.
+3.
+To assess robustness to external perturbations, we conducted experiments in which the quadrotor was required to track these trajectories while being subjected to the disturbances $w_{x},w_{y},w_{z}$ were randomly sampled from the interval [0.2,0.2][-0.2,0.2].
+Each algorithm was evaluated over 500 episodes and assessed using the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})$ over a single episode.
 6.2 Results
-Figure 3: Mean and standard deviation of the total episode cost.
-RDDPG achieves the lowest average cost and the smallest variance across all scenarios.
-This indicates that the RDDPG yields a robust agent that performs consistently even under severe external disturbances.
-[…]
+Figure 2 compares the performance of RDDPG with TD3, RARL (RARL), PR-MDP, and NR-MDP  (AR-MDP) across two evaluation trajectories under random disturbances.
+Across both trajectories experiments, RDDPG consistently demonstrates the most stable and reliable performance.
+Its cost distribution exhibits the smallest variance among all evaluated methods, indicating that the learned policy maintains robust tracking behavior even under random disturbances.
+In contrast, TD3 and RARL often achieve higher peak rewards but suffer from significantly larger variance, revealing high sensitivity to perturbations and unstable control behaviors.
+PR-MDP and NR-MDP show more moderate performance with reduced variability compared to TD3 and RARL, yet they still fall short of the consistency exhibited by RDDPG.
 By maintaining a consistently low and stable cost across a wide range of disturbance intensities, RDDPG exhibits superior robustness and reliable control performance, even under challenging environmental conditions.
@@
 By optimizing the policy against worst-case disturbances during training, RDDPG learns disturbance-resilient strategies that improve generalization and ensure stable performance under unseen or highly uncertain conditions.
-Figure 4: Trajectory results of all algorithms under the $\pm 15$ wind scenario and Traj (1).
-Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds.
-RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms.
-The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure 5.
-Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC.
-In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios.
-However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG.
-Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes.
-Detailed numerical results are provided in at Appendix Table 5.
-Figure 5: Mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each time step.
-RDDPG consistently achieves low error and standard deviation across all scenarios, whereas other methods exhibit unstable performance.
 7 Conclusion
@@
 In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances.
-Furthermore, we introduce RDDOG, which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks.
+Furthermore, we introduce RDDPG, which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks.
 In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances.
 Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios.
-These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.
+These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability.
 Appendix A Hyperparameters Used in Training
@@
 $\epsilon$ Coefficient for angle error 11
-$\lambda$ Living reward 10-10
 $\zeta$ Penalty $1000$
-A.3 DDPG and TD3
-Table 3: Caption
+A.3 Algorithm parameters
+Table 3: Algorithm parameters
 Parameter Description Value
@@
 $\sigma$ Exploration noise 0.2
-A.4 SAC
-Table 4: SAC hyper-parameters
-Parameter Description Value
-$\alpha_{actor}$ Learning rate for user 0.0003
-$\alpha_{critic}$ Learning rate for critic 0.003
-$|B|$ Batch size 256
-$\gamma$ Discounted factor 0.99
-$\alpha_{entropy}$ Entropy coefficient 0.99
-Appendix B Experiment results
-Table 5: The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}$ between the target and the quadrotor. Bold indicates the best performance algorithm
-$\pm 5$ wind speed $\pm 10$ wind speed $\pm 15$ wind speed
-Traj (1) Traj (2) Traj (1) Traj (2) Traj (1) Traj (2)
-[…]
-[…]
-[…]
-[…]
-Figure 6: In $\pm 15$ m/s wind scenario
-Figure 7: In $\pm 15$ m/s wind scenario
-Figure 8: In $\pm 15$ m/s wind scenario
-Figure 9: In $\pm 15$ m/s wind scenario
-Figure 10: In $\pm 15$ m/s wind scenario
+$\eta$ Coefficient of disturbance norm 10
+All algorithms were trained using the same set of hyperparameters shown in the table.
+For RDDPG, the scalar coefficient $\eta$ multiplying the disturbance norm was fixed to $10$.
+For PR-MDP and NR-MDP, the adversarial noise injection probability $\alpha_{\mathrm{AR}}$ was set to $0.1$.